### v3

# Strategic Tradeoffs Between Humans and AI in Multi-Agent Bargaining

Abstract
As large language models (LLMs) are increasingly embedded in collaborative human activities such as business negotiations and group coordination, it becomes critical to evaluate both the performance gains they can achieve and how they interact in dynamic, multi-agent environments. Unlike traditional statistical agents such as Bayesian models, which may excel under well-specified conditions, large language models (LLMs) can generalize across diverse, real-world scenarios, raising new questions about how their strategies and behaviors compare to those of humans and other agent types. In this work, we compare outcomes and behavioral dynamics across humans (N = 216), LLMs (GPT-4o, Gemini 1.5 Pro), and Bayesian agents in a dynamic negotiation setting under identical conditions. Bayesian agents extract the highest surplus through aggressive optimization, at the cost of frequent trade rejections. Humans and LLMs achieve similar overall surplus, but through distinct behaviors: LLMs favor conservative, concessionary trades with few rejections, while humans employ more strategic, risk-taking, and fairness-oriented behaviors. Thus, we find that performance parity—a common benchmark in agent evaluation—can conceal fundamental differences in process and alignment, which are critical for practical deployment in real-world coordination tasks. By establishing foundational behavioral baselines under matched conditions, this work provides a baseline for future studies in more applied, variable-rich environments.
1 Introduction
Historically, bespoke statistical models have excelled in structured economic exchanges with well-defined action spaces (Sun and Müller, 2013; Dehghanpour et al., 2016). More recently, the generalizability and flexibility of large language models (LLMs) have enabled delegation in unstructured contexts: negotiating supply chain contracts (Van Hoek et al., 2022), coordinating international trading (Corvin, 2024), and facilitating online commerce transactions (PYMNTS, 2025; Gaarlandt et al., 2025). Unlike more structured models, LLMs can generalize under uncertainty and partial information, reducing the need for task-specific customization. As AI agents grow more sophisticated and are entrusted with higher-stakes decisions, understanding how different types of agents behave in dynamic, multi-agent settings is essential. In particular, navigating the trade-offs between agent types—not only in terms of negotiated outcomes, but also in their alignment with human strategies and norms—is critical for responsible delegation.
A barrier to understanding delegation tradeoffs has been the difficulty of evaluating disparate agent types in comparable interactive settings. Static QA-style benchmarks fail to capture the dynamic, multi-agent interactions that characterize real-world negotiation. Indeed, recent work has called for more realistic evaluation environments that reflect the social and strategic complexities faced by deployed AI agents (Raman et al., 2024; Goktas et al., 2025).
We address this challenge by introducing a novel multi-player bargaining game that enables direct comparison of different agent types (e.g., humans, LLMs, and statistical models) under identical conditions. In our game, players receive endowments of colored chips with private valuations and negotiate trades to maximize their individual surplus. Bargaining is an ideal testbed for dynamic, multi-agent evaluations, as it inherently exposes tensions between social norms, individual preferences, and strategic reasoning, reflecting the complexities of real-world environments. Moreover, from a theoretical perspective, bargaining under partial information admits no Pareto efficient, incentive-compatible solutions (Myerson and Satterthwaite, 1983a), forcing agents to navigate trade-offs between individual gain and mutually beneficial cooperation.
We conduct an empirical study of our game across three types of agents: (i) human participants (N = 216), (ii) LLM-based agents, and (iii) specialized Bayesian agents, all facing identical randomized endowments. The Bayesian agents achieved the highest surplus, proposing aggressive bargains at the cost of frequent rejections. LLMs matched human surplus in certain game conditions but exhibited fundamentally different behaviors: where humans made fairness-driven concessions, LLMs favored conservative, low-variance trades with high acceptance rates.
More broadly, these findings expose a limitation of outcome-based evaluation: aggregate surplus can mask important differences in behavior and alignment. As AI systems increasingly mediate real-world negotiations, it becomes essential to assess not just what agents achieve, but how they navigate trade-offs, adjust to strategic dynamics, and engage with social norms. The primary contributions of this work are threefold:
  • A dynamic multi-agent environment for comparative evaluation. We design and implement a bargaining game that enables direct, identical-condition comparisons of negotiation strategies across different agent types, with game variations of increasing complexity. We accompany the platform with a linear programming formulation of the Pareto-optimal surplus, providing a theoretical benchmark for performance.
  • An empirical characterization of human, LLM, and Bayesian negotiation strategies. Through a study with 216 human participants (N=216), we reveal distinct behavioral profiles: Bayesian agents adopt aggressive, surplus-maximizing strategies with high rejection rates; humans exhibit strategic concessions and fairness norms; LLMs act as conservative agents that yield high rates of offer acceptance.
  • A framework to demonstrate how performance parity can conceal strategic misalignment. Our experimental benchmarks show that aggregate outcomes (i.e., total surplus) are insufficient for evaluation, as they mask fundamental differences in agent behavior. This finding highlights the need to move beyond performance-based metrics to assess process and alignment for the responsible deployment of agents in real-world coordination tasks.
2 Related work
Statistical agents have excelled in well-defined tasks with explicit objectives, preferences, and constrained action spaces (Sun and Müller, 2013; Dehghanpour et al., 2016). Recent advances in LLMs have broadened the scope of delegation to include tasks requiring broader social reasoning, such as preference elicitation, strategic pricing, natural-language negotiation, and alliance formation in multi-agent games like Diplomacy (Fish et al., 2024; Soumalias et al., 2025; Tessler et al., 2024; Vaccaro et al., 2025; Lubars and Tan, 2019; Bakhtin et al., 2022; Kramár et al., 2022; Deng et al., 2024). LLMs have exhibited emergent capabilities in economic computation and human alignment, replicating behaviors in bargaining, auctions, and other economic games (Horton, 2023; Aher et al., 2022; Argyle et al., 2022), and reflecting human subgroup-specific patterns when conditioned on demographic traits (Manning et al., 2024; Xie et al., 2024; Zhu et al., 2024; Park et al., 2023; Qian et al., 2025).
However, appropriate deployment and delegation depend not only on system capabilities (Jahani et al., 2025), but also on user expectations, skills, and alignment with social norms; users may suboptimally delegate when system behavior violates expectations (Mozannar et al., 2023; Agarwal et al., 2023; Mullainathan and Obermeyer, 2021; Palminteri et al., 2024; Kapania et al., 2022; Qian and Wexler, 2024). Addressing alignment gaps requires evaluating both capabilities and behaviors in dynamic, social contexts—a need recognized by recent calls for rigorous methods to steer and assess agent behavior in strategic multi-agent settings (Raman et al., 2024; Goktas et al., 2025). To concretely examine these challenges, we study how humans, LLM-based agents, and Bayesian models behave in a dynamic bargaining task, offering a comparative lens on alignment in strategic social interactions.
3 The bargaining game
3.1 Motivation from real-world negotiations
Negotiation remains one of humanity’s most fundamental coordination tasks Nash et al. (1950); Rubinstein (1982). Bilateral and multi-party bargaining appear in contexts such as three-team NBA tradesPelechrinis et al. (2023), B2B inventory swaps, and corporate barter networks Binmore (2007). These environments require participants to negotiate over bundles of heterogeneous assets with private valuations, making offers such as “X of A for Y of B (plus cash),” under real-world feasibility and time constraints.
Such bargaining is both strategically and communicatively demanding Farrell and Rabin (1996); Camerer (2003), motivating efforts to automate or augment the process. With advances in large language models (LLMs), it is now plausible to automate not just basic proposals, but also reasoning, persuasion, and adaptation in natural language. However, the strategic differences across LLM and human negotiation behavior remain underexplored; LLMs may diverge from human norms in ways that shape outcomes, affecting agreement rates, surplus allocation, and perceptions of fairness.
Studying these trade-offs in applied settings invites confounds; for example, if we were to evaluate the exchange of real commodities like secondhand goods in an online marketplace, participants’ heterogeneous preferences would confound direct comparison across agent types. To address this, we construct an abstract bargaining game that captures the essential structure of bargaining markets: fixed assets, a shared numeraire, private valuations, and an offer–response protocol Roth and Kagel (1995). The tensions embedded in our game—private information, strategic risk, and the trade-off between individual gain and mutual benefit—are fundamental to a wide range of real-world coordination tasks, from collective bargaining for labor contracts to negotiating international climate treaties Schelling (1980); Myerson and Satterthwaite (1983b). Therefore, our findings on agent behavior provide a foundational and generalizable diagnostic for understanding how these agents will likely act in more complex, unstructured interactions.
3.2 A stylized game
We introduce a multi-player bargaining game which captures strategic trade-offs, social signaling, and coordination challenges, influenced by real-world negotiations. Figure 1 presents an overview of the game, in which three players aim to maximize individual surplus over nine turns. Each begins with ten chips of each color. Green chips act as a shared numeraire valued at \$0.50 by all; other chip colors (red, blue, purple) have private values, drawn uniformly between \$0.10 and \$1.00. Game complexity is varied by incrementally adding chip types: red, then blue, then purple, with green always present. [^1] We show in Appendix A.2 that decision complexity grows with the number of chip colors.
Figure 1: This provides a broad overview of the bargaining game. Left: Game setup with initial chip counts and valuations. Center: Simplified version of gameplay (offer proposals and acceptances). Right: The types of agents empirically evaluated and simulated in this study.
Gameplay.
In a fixed, random order, players propose trades by offering chip(s) of one color in exchange for another. The two non-proposers simultaneously decide whether to accept or decline the proposal; if both accept, one is chosen at random to clear the transaction. Players can only propose and accept offers that they have sufficient inventory to fulfill. All players observe transactions and chip holdings, but not others’ chip valuations. At the end of the game, each player receives a payout based on the surplus generated from their final chip holdings.
4 A pareto optimal upper bound
Before exploring empirical outcomes, we provide a theoretical upper bound analysis. Although our bargaining game lacks a tractable exact game-theoretic solution, we can compute a Pareto-optimal allocation via linear programming to benchmark agent performance in our empirical results.
4.1 Notation
Consider a game M = (I, G, v, a) with a set of agents I and chips G. Each agent $i$ has a personal valuation $v_{ig}$ for each good $g$. These valuations are static, independent across agents, and complete for each agent-good pair. All agents have an initial allocation $a^{0}_{ig}$ of each chip, which indicates the amount of chip $g$ allocated to agent $i$. Given these allocations, an agent’s welfare $w_{i}$ is the sum of their valuations weighted by their allocations: $w_{i}=\sum_{g\in G}v_{ig}a_{ig}$. The total welfare w is thus the sum of all individual welfare: $\textbf{w}=\sum_{i\in I}w_{i}=\sum_{i\in I}\sum_{g\in G}v_{ig}a_{ig}$. We will refer to the difference between initial and post-trade welfare as surplus gain in the following analysis.
4.2 Efficient allocations
In our game, agents trade chips to improve their welfare. Given their initial endowments $a^{0}_{ig}$, mutually beneficial trades may exist that increase total welfare. We provide an upper bound to the obtainable total welfare, assuming the Pareto condition that no player ends with worse utility than they began.
Definition 1.
(Optimal Allocation of Chips). An optimal allocation of chips is any set of allocations $\mathbf{A^{*}}=\{a^{*}_{ig}\ |\ i\in I,g\in G\}$ that solves:
$\operatorname*{\arg\!\max}_{\mathbf{A}}\sum_{i\in I}\sum_{g\in G}v_{ig}a_{ig}$
subject to:
  1. (i)
    Conservation of goods: $\sum_{i\in I}a_{ig}=\sum_{i\in I}a^{0}_{ig},\forall\ g\in G$
  2. (ii)
    Pareto improvement: $\sum_{g\in G}v_{ig}a_{ig}\ \geq\ \sum_{g\in G}v_{ig}a^{0}_{ig},\forall\ i\in I$.
  3. (iii)
    Non-negativity: $a_{ig}\geq 0\forall\ i\in I,g\in G$.
Intuitively, in a centralized setting with full information, $\mathbf{A^{*}}$ represents a Pareto-efficient allocation that maximizes aggregate utility without making any agent worse off. While this optimal allocation $\mathbf{A^{*}}$ may not be unique, the maximum total welfare $\mathbf{w^{*}}$ achieved at the optimum is unique. This optimization problem can be solved efficiently using interior point methods. Its computational complexity is bounded by $O\bigl((|I||G|)^{3.5}\bigr)$. Note that this upper bound is unlikely to be achieved in practice, as no mechanism for decentralized bargaining with private information can guarantee both Pareto efficiency and incentive compatibility Myerson and Satterthwaite (1983a). [^2] A proof that players of the bargaining game lack a dominant strategy is provided in AppendixA.1.
5 Experimental setup
In our design, human participants first completed games in groups of three. Each group was randomly assigned to either the 2-chip, 3-chip, or 4-chip game variation, allowing us to evaluate performance and strategic trade-offs under increasing task complexity. We then replicated each game (identical initial chip endowments) with LLM- and Bayesian- agent groups, enabling direct comparison across the groups. [^3] For reproducibility, details on the human game interface are provided in Appendix B.1. Full LLM prompts are provided in Appendix G, and the Bayesian agent algorithms are detailed in Appendix A.3.
5.1 Human subjects
We recruited 216 U.S.-based participants from Prolific to complete the bargaining game on Deliberate Lab (Tsai et al., 2024). [^4] IRB approval and informed consent were obtained; we applied no inclusion criteria for recruitment. Participants completed two games over a 30-minute session, which included time waiting for two others to join a live game. To incentivize strategic play, participants received the surplus value of their final chip holdings in addition to a \$4 base payment. The average total payout was \$12.24 $\pm$ \$6.12. Each participant completed two games with different chip valuations and profiles. For each of the 2-, 3-, and 4-chip game variants, 24 groups of three players played 2 games each, yielding $N=3\times 24\times 3\times 2=432$ effective participants across 144 total games. [^5] The two games are evaluated independently, as ordering and learning effects were minimal.
5.2 LLM agents
We then replicated each human game with LLM agents. At each turn, an LLM agent received the same information as its human counterpart, including private chip valuations, public chip holdings, and trade history. One agent proposed a trade per round; the others evaluated it.
We developed two types of agents to evaluate the tradeoffs between simple prompting and chain-of-thought refinement. As our primary goal was to evaluate baseline performance under minimal optimization, we first constructed an out-of-box agent used simple prompts adapted from the human instructions. To assess the benefit of lightweight refinement, we also implemented a refined agent which generated candidate proposals and selected one using “type-2” reasoning, mimicking slower, more deliberate decision-making (Furniturewala et al., 2024). Each agent type was implemented across two model families for generalizability: OpenAI’s GPT-4o (OpenAI et al., 2024) and Google’s Gemini 1.5 Pro (Team et al., 2024), which were the latest full-capacity models from each family at time of writing. [^6] For robustness, we also benchmarked the latest lightweight models (Gemini 2.5 Flash, GPT-o4-mini), but they failed to execute basic surplus-maximizing trades. We therefore restrict our main analysis to full-capacity models; see Appendix F for additional context. Prompt temperature was set to 0.5, following mid-range recommendations for structured reasoning tasks that require coherence and variability (Arora et al., 2024). Prompts were scaffolded using a chain-of-thought strategy through the open-source EDSL library (Wei et al., 2022; Horton et al., 2024).
5.3 Bayesian agents
We include Bayesian players as a statistical benchmark: these agents represent the traditional, hard-coded approach to negotiation, explicitly programmed to optimize within the structure of this bargaining game. Each agent i forms a joint belief about the valuations of all other players, $\mathbf{v}_{-i}$. When agent i proposes a trade $(x_{g},y_{r})$ – offering $x$ chips of type $g$ in exchange for $y$ chips of type $r$ – they choose a trade to maximize their expected payoff:
$\max_{(\mathbf{x},\mathbf{y})}\sum_{\mathbf{v}_{-i}}\mathbf{1}\Bigl\{\exists\,j\neq i:\text{accept}\bigl(v_{j},x_{g},y_{r}\bigr)\Bigr\}\;\times\;\Delta u_{i}(v_{i},x_{g},y_{r})\;\times\;B_{i}\bigl(\mathbf{v}_{-i}\bigr),$ (1)
where $\text{accept}\bigl(v_{j},x_{g},y_{r}\bigr)$ indicates that player j with valuation $v_{j}$ would accept the trade, $\Delta u_{i}(v_{i},x_{g},y_{r})=v_{i,r}y_{r}-v_{i,g}x_{g}$ is player $i$’s change in utility and $B_{i}(\mathbf{v}_{-i})$ is player $i$’s belief (probability) that the other players’ valuations are $\mathbf{v}_{-i}$. This agent assumes that their opponent’s acceptance decision is myopically rational, i.e. a receiver with valuation $v_{j}$ accepts if and only if they have sufficient chips for the trade and their expected utility is positive, i.e. $v_{j,g}x_{g}-v_{j,r}y_{r}>0$.
All players update their priors after observing acceptance decisions. A player discards valuation states $\{v_{j}\}$ that contradict the observed event, assuming myopically rational opponents, and re-normalizes the remaining probabilities. When the trade is accepted, the proposer keeps only valuations that would be accepted in its distribution over the chosen receiver; the receiver keeps only valuations that would be accepted in its distribution over the proposer. Bystanders update priors for both accordingly. When there is no trade, all players discard valuations that conflict with the observed non-acceptance.
Figure 2: Surplus trajectories for human, LLM, and Bayesian-learning agent simulations across game complexities. Each unique game yields a blue line corresponding to the ratio of surplus achieved relative to the computed optimal allocation. Means and 95% confidence intervals are highlighted and listed in Appendix 3. All populations generated positive aggregate surplus.
6 Results
6.1 Performance
Figure 2 shows surplus trajectories by agent type and game variation. Bayesian agents reached 74% of the Pareto-efficient optimum in the two-chip game (upper right panel). Across all variations, Bayesian agents achieved the highest surplus. GPT-4o performed comparably to human participants; however, humans statistically outperformed Gemini 1.5 Pro. [^7] This may reflect the recency of such capabilities; Gemini 1.5 Pro was released three months prior to GPT-4o and represented an earlier generation of model development. Refined LLMs did not significantly impact surplus compared to out-of-the-box LLMs (Appendix E).
6.2 Procedural alignment
In addition to performance, we analyze procedural alignment, how the agents engage in the mechanics of negotiation, through two lenses: 1) observed trading patterns, and 2) regret minimization strategies.
6.2.1 Trading patterns and volume dynamics
Figure 3: Trading patterns in the 3-chip game, visualizing (i) net surplus change for the proposer and (ii) trade ratio. Accepted trades are shown in green, rejected trades in red. Marginal distributions across each axis are adjacent to the plots. The vertical dashed line marks zero net surplus (no net value created), and the horizontal solid line marks balanced exchange (1:1 ratio of chips). Figure values are provided in Table 5.
Humans and social norms.
Figure 3 plots individual trade proposals for the three-chip game for each agent type. Results for the two- and four-chip games are qualitatively similar (Appendix F.2). The vertical balanced value line indicates proposals that generated no net surplus, while the horizontal balanced ratio line corresponds to 1:1 exchanges of chips. In the left panel, we see that the human proposals cluster around the balanced ratio line: that is, humans often offer a similar number of chips as they are requesting. Most trade proposals lie to the right of the balanced value line (i.e., provide positive surplus to the proposer), though some reduce total surplus.
LLMs and concessionary exchanges.
Both LLM families propose trades near the vertical balanced net value line, yielding little total surplus with no anchoring to balanced trade ratios. A noticeable vertical tail of proposals extends above the 1:1 balanced ratio line, indicating a tendency to offer more chips than requested (even up to 5:1). Many proposals also incur a net surplus loss.
Bayesian agents and rational surplus maximization.
By construction, Bayesian agents strictly maximize surplus (evidenced by positive surplus proposals) and avoid net-loss trades (no points left of the balanced value line). Rather than the 1:1 balanced ratio exhibited in human proposals, the Bayesian offers cluster below the parity line (ratio < 1), meaning they ask for more chips than they give. Proposal rejections are high.
Trade pattern comparisons.
In sum, these trading patterns reveal clear strategic differences: humans offer trades suggesting a consideration for balanced, fair exchanges (Danz et al., 2022), LLMs adopt a “concessionary posture” with high proposal acceptance rates, and Bayesian agents adopt a value-extractive strategy (Fisher et al., 2011), asking to receive more than they give and incurring higher rejection rates. To evaluate whether any of these strategies were more optimal ex-post, we next evaluate these trades using a regret-minimization framework.
6.2.2 Strategic alignment and regret minimization
At each turn, a participant might take an action of proposing, accepting, or declining a trade. In no regret actions, no feasible later alternative would yield a better valuation for the target chip. Forced regret occurs when a higher-surplus option arises later, but the agent’s earlier decisions—such as premature trades or overcommitting inventory—prevent acting on it. Finally, unforced regret captures scenarios where a better option becomes available, yet the agent fails to take it.
To avoid speculative modeling of unobserved strategies, we limit the scope of our analysis. First, we only consider myopically rational transactions (i.e., positive surplus) for proposers, acceptors, or decliners. [^8] Bayesian agents satisfy this by design. Interpreting irrational trades, which may be driven reciprocity, signaling, or cognitive error, would require assumptions we cannot credibly test. Second, when evaluating proposer behavior, we focus on accepted trades; rejected offers may result from strategic signaling by either party or proposer misjudgment of demand. [^9] This analysis is distinct from evaluations conducted from the perspective of decliners. Third, “acceptors” include those who intended to accept a trade, not only those who were selected to trade. [^10] This occurs when both accept the trade and one is randomly chosen. This preserves the behavioral signal of willingness to accept, without conflating risk tolerance or random assignment.
Figure 4 shows an example of a player’s valuations as they trade over the course of a single game. [^11] The sum of all players’ surplus changes at each turn correspond to the blue lines in Figure 2. On turn 1, Player 3 exhibited a forced regret action as an acceptor: they accepted a trade which resulted in their having insufficient chips to accept a more profitable offer in turn 2. On turn 7, Player 2 exhibited an unforced regret action as a decliner; they declined a profitable trade and did not recover the profits later. On Turn 8, Player 2 made a no regret proposal; this was the best deal they could have made, as there were no better future trades. [^12] More detailed explanations are provided in Appendix D.
Figure 4: Example of player chip value trajectories over nine turns of a trading game (3-Chip). To avoid speculation, we do not analyze analyze Player 1’s proposal on Turn 4 (an unaccepted trade), or Player 3’s acceptance on Turn 5 (accepting a negative surplus offer).
Figure 5: Counts of strategic actions across games. Bayesian agents performed more optimal, no-regret actions relative to other populations as game complexity increased. Declining a proposal would not cause forced regret, as this action does not commit any chips. The total number of actions naturally increases with game complexity, reflecting the larger solution space, not necessarily improved decision-making.
Figure 5 shows the classifications of actions across games and populations. [^13] We exclude refined LLMs as they behave similarly to out-of-the-box models; see Appendix F.1. The relative performance of Bayesian agents improves as game complexity increases. Combined with the surplus outcomes in Figure 2, this suggests that Bayesian agents are not necessarily improving, but rather that LLM and human performance degrades as the action space expands.
Bayesian agents generate a higher proportion of optimal (no regret) proposals and fewer regrettable ones compared to LLMs and humans, reflecting more effective surplus-maximizing calculations. However, Bayesian agents also register the most declines—both regrettable and non-regrettable. This indicates that their proposal strategies prioritize their own advantage, without necessarily considering whether those offers will be accepted by others.
In contrast, humans and LLMs exhibit more regrettable proposals and acceptances relative to Bayesian agents. They also exhibit fewer declines overall, indicating that they propose more agreeable offers at the cost of making less optimal actions. Differences in forced regret actions are relatively modest, though LLMs show slightly higher forced regret counts.
7 Discussion
Humans reflect economic and social reasoning.
Humans offered trades consistent with classical fairness norms, proposing a balanced trade ratio. While these proposals were less frequently optimal than that of Bayesian agents, they were also accepted more often. Post-game survey responses consistently emphasized fairness and cooperation as decision drivers, suggesting that humans prioritize social norms, even in one-shot interactions where long-term reputation effects would not affect payoff. [^14] Post-game survey questions and quantitative responses are in Appendix C. However, nuanced human social motivations are difficult to formalize into algorithms, posing a fundamental challenge when delegating coordination and negotiation to artificial agents.
LLMs can perform on-par, but may be concessionary by design.
The latest model, GPT-4o, achieved human-level surplus with minimal prompting, demonstrating flexibility when applied to novel tasks. However, their behavior was compromising and reactionary: they favored value-balanced proposals which were less optimal than Bayesian proposals but more likely to be accepted, and exhibited high rates of forced regret trades, suggesting limited strategic foresight in dynamic scenarios. This conservative play style may arise from several factors: (i) training on cooperative, information-sharing dialogues, where minimizing friction is implicitly rewarded (Wei et al., 2022); (ii) a tendency to generate risk-averse, passive responses compounded by a lack of outcome-driven feedback (Ouyang et al., 2022); and/or (iii) a possible learned aversion to asymmetric outcomes, where generous offers serve as a coordination strategy to secure agreement (Fisher et al., 2011). In adversarial or zero-sum negotiations, this overly concessionary tendency could lead LLMs to systematically overcompromise.
Bayesian agents outperform, but lack social adaptability.
Bayesian agents excelled at surplus-maximizing proposals, aggressively extracting value. This strong performance was constructed through a hand-crafted algorithm specifically aligned to the game’s incentives. In contrast, humans and LLMs relied on their flexible “out of the box” capabilities, lacking such tailored inductive biases.
In practice, Bayesian agents’ extractive, rejection-tolerant strategy might fail in real-world negotiations, especially in repeated interactions requiring trust and reciprocity. Extending Bayesian agents to more socially compatible trading would require richer models of preferences, fairness, and adaptive behavior, which could increase complexity and decrease robustness.
Hybrid models offer promise, but need social reasoning.
We find that humans consider social factors, LLMs favor safe, consensus-seeking strategies, and Bayesian agents excel through narrow, task-specific optimization. These contrasting strengths suggest potential for complementary approaches, where LLMs are augmented with Bayesian tools or planning modules to improve foresight, such as inventory management and lookahead reasoning. However, negotiation is not just a planning problem; success also depends on modeling others’ intentions, social norms, and the unwritten rules of cooperation. Without explicit social reasoning, even hybrid agents will struggle with the social nuances that humans navigate instinctively.
Procedural alignment matters for human-AI interaction.
Furthermore, how agents negotiate is as important as what they achieve. Despite humans and LLMs yielding similar aggregate surplus in some contexts, they exhibited distinct strategies. Crucially, these differences only became visible through a custom experimental design that enabled matched, one-to-one comparisons across populations; simple outcome-level aggregation would have obscured these divergences. As AI agents increasingly participate in human-facing decision-making, focusing solely on efficiency metrics risks missing misalignments in process, intent, and social compatibility.
Efficiency and scale advantages.
A final consideration beyond strategy is scale. LLMs and Bayesian models can resolve negotiations near-instantaneously, unaffected by cognitive load or decision fatigue. LLMs, in particular, are faster to deploy out-of-the-box compared to bespoke Bayesian models; in our case, our out-of-the-box agents required minimal prompt scaffolding around the human instructions, whereas we constructed a custom algorithm for the Bayesian agents. These efficiency gains make automation attractive for high-volume or time-sensitive negotiations, even if strategic or social limitations remain.
Social impacts and risks.
Agents that achieve surplus-yielding outcomes through fundamentally different negotiation styles pose risks for trust, fairness, and social acceptance. If deployed in real-world negotiations, models like LLMs may default to safe, concessionary behavior, leading to suboptimal deals and reinforcing status-quo power dynamics. Conversely, optimization-focused agents like Bayesians may extract disproportionate value, undermining perceptions of fairness and cooperation. These dynamics may lead to unexpected interaction effects in mixed-agent environments. Designing AI agents that are not only strategically capable but also socially compatible will be critical for their long-term acceptance and responsible deployment.
8 Limitations and future work
Our study focuses on a single, stylized negotiation game with static valuations and single shot interactions, conditions that align well with Bayesian agents’ task-specific optimization and may exaggerate their advantage. While humans and LLMs bring general-purpose reasoning to the task, their broader capabilities in dynamic, multi-round, or reputation-driven negotiations remain unexplored. Furthermore, our LLMs used minimal prompting, reflecting realistic deployment but limiting exploration of more strategic behaviors from alternative prompts or fine-tuning.
Beyond performance metrics, our analysis reveals distinct behavioral profiles but does not disentangle whether these arise from reasoning limitations, social norms, or inductive biases inherent to each system. Future work should investigate how humans perceive and trust AI negotiators with different procedural styles. Another promising avenue for future work is to allow agents to communicate with natural language. Finally, developing methods to steer LLM agents towards desired procedural norms explicitly represents a significant ongoing challenge and opportunity.
9 Conclusion
This study demonstrates an empirical approach for comparing agents in dynamic, social settings. By examining human, LLM, and Bayesian behavior in a structured bargaining game, we show how performance measures can conceal fundamentally different strategic approaches. As AI systems are increasingly delegated complex tasks, understanding how their decision-making processes align with human expectations and social values is critical to fostering trust, effective cooperation, and responsible deployment in real-world applications.
Appendix A Algorithms and proofs
A.1 Proof of no dominant strategy
Claim.
No agent $i$ $\in$ $\mathbf{I}$ in the game $\mathcal{M}$ has a dominant strategy.
Setup.
Let $\mathbf{I}$ be the set of agents, $\mathbf{G}$ be the set of goods. For each agent $i\in\mathbf{I}$ and good $g\in\mathbf{G}$, there is a valuation $v_{ig}\in\mathbb{R}_{\geq 0}$. The (feasible) allocations are given by $a_{ig}$, the amount of good $g$ allocated to agent $i$.
The welfare of agent $i$ is
$w_{i}\;=\;\sum_{g\in\mathbf{G}}\,v_{ig}\,a_{ig}.$
The total social welfare is
$w\;=\;\sum_{i\in\mathbf{I}}\,w_{i}\;=\;\sum_{i\in\mathbf{I}}\sum_{g\in\mathbf{G}}v_{ig}\,a_{ig}.$
Let each agent $i$ have a strategy space $\mathcal{S}_{i}$. A (pure) strategy $s_{i}\in\mathcal{S}_{i}$ is a complete description of agent $i$’s behavior in the market (e.g., how they propose, reject, accept, etc.). We write $\mathbf{s}_{-i}$ for a strategy profile of all agents other than $i$. Let $u_{i}(s_{i},\mathbf{s}_{-i})$ denote agent $i$’s resulting payoff (or final utility/welfare) under strategy $s_{i}$ against opponents’ strategies $\mathbf{s}_{-i}$. A strategy $s_{i}^{*}$ is called dominant for agent $i$ if, for all $\mathbf{s}_{-i}\in\mathcal{S}_{-i}$,
$u_{i}\bigl(s_{i}^{*},\mathbf{s}_{-i}\bigr)\;\;\geq\;\;u_{i}\bigl(s_{i},\mathbf{s}_{-i}\bigr)\quad\text{for all }s_{i}\in\mathcal{S}_{i}.$
In other words, $s_{i}^{*}$ guarantees the highest possible utility for $i$, irrespective of what other players do.
Proof.
Assume, by way of contradiction, that there exists a dominant strategy $s_{i}^{*}$ for some agent $i$. By the definition of dominance, $s_{i}^{*}$ must satisfy
$u_{i}\bigl(s_{i}^{*},\mathbf{s}_{-i}\bigr)\;\geq\;u_{i}\bigl(s_{i},\mathbf{s}_{-i}\bigr)\quad\text{for all }s_{i}\in\mathcal{S}_{i}\text{ and for all }\mathbf{s}_{-i}\in\mathcal{S}_{-i}.$
Construct two different scenarios of other agents’ strategies and valuations, denoted by $\mathbf{s}_{-i}^{A}$ and $\mathbf{s}_{-i}^{B}$. These scenarios can differ in 1) how other agents value the goods $\{v_{jg}:j\neq i\}$, and 2) the demands or offers they make (i.e. how $\mathbf{s}_{-i}$ translates into allocations $a_{jg}$).
  • In scenario $A$, suppose the other agents are willing to trade or allocate a certain good $g^{*}$ to agent $i$ only if $i$ offers a high price or concedes in some manner. A specialized alternative strategy $s_{i}^{A}$ (instead of $s_{i}^{*}$) might yield a strictly higher payoff if $i$ trades aggressively for $g^{*}$.
  • In scenario $B$, suppose the others behave differently (e.g., they no longer place much value on $g^{*}$ but highly value some other good). Now a different specialized strategy $s_{i}^{B}$ might yield a strictly higher payoff (because trading strategy for $g^{*}$ in scenario $A$ is no longer beneficial here).
Contradiction of dominance: Because scenarios $A$ and $B$ are both feasible profiles of opponents’ behavior ($\mathbf{s}_{-i}^{A}$ and $\mathbf{s}_{-i}^{B}$), there is no single strategy $s_{i}^{*}$ that can simultaneously guarantee at least as high a payoff as both $s_{i}^{A}$ and $s_{i}^{B}$ across those distinct opponent profiles. Formally,
$\exists\,\mathbf{s}_{-i}^{A}:u_{i}\bigl(s_{i}^{A},\,\mathbf{s}_{-i}^{A}\bigr)\;>\;u_{i}\bigl(s_{i}^{*},\,\mathbf{s}_{-i}^{A}\bigr),\quad\text{and}\quad\exists\,\mathbf{s}_{-i}^{B}:u_{i}\bigl(s_{i}^{B},\,\mathbf{s}_{-i}^{B}\bigr)\;>\;u_{i}\bigl(s_{i}^{*},\,\mathbf{s}_{-i}^{B}\bigr).$
This contradicts the definition of $s_{i}^{*}$ as a dominant strategy.
Because the above argument does not rely on any special assumption beyond the existence of at least two different plausible profiles $\mathbf{s}_{-i}^{A}$ and $\mathbf{s}_{-i}^{B}$, we conclude that no strategy for agent $i$ can be dominant. Since $i$ was arbitrary, no agent in the game has a dominant strategy. $\quad\square$
A.2 Empirical computation of increasing game complexity
Claim.
The decision space increases quadratically as the number of chip colors increases.
Proof.
This is a simple consequence that in games with $k$ different colors of chips, there are $\binom{k}{2}$ possible pairs of chips to transact. Since the distribution of opponents’ valuations is symmetric across chips (except the green numeraire chip), each proposer must consider $O(k^{2})$ different feasible trades. Indeed, if we empirically compute the expected number of myopically rational trades [^15] We define a myopically rational trade to be one where the proposer gains surplus and assigns some probability that at least one opponent would gain surplus by accepting the trade. in the starting configuration of the 2, 3, and 4-chip game, we arrive at 37.1, 120.7, 250.7 respectively. That is the 3-chip game has approximately $\binom{3}{2}$ as many myopically rational trades in expectation as the 2-chip game, and the 4-chip game has approximately $\binom{4}{2}$. This approximation becomes more exact as the number of chips increases due to the reduced relative impact of the numeraire.
A.3 Multi-agent trading algorithm with Bayesian learning
Algorithm 1 Multi-agent Trading Algorithm with Bayesian Learning
1:Inputs:
[…]
7:for all $i\in\mathbf{I}$ do
8:  initialize $B_{i}(\mathbf{v}_{-i})$ $\triangleright$ Each agent’s prior over other agents’ valuations
9:end for
[…]
11:while possible_tradespossible\_trades do
12:  $trades\_found\leftarrow\texttt{false}$
13:  $I_{\text{shuffled}}\leftarrow\text{RandomShuffle}(\mathbf{I})$
14:  for $k=1$ to $|\mathbf{I_{\text{shuffled}}}|$ do
15:   for $l=k+1$ to $|\mathbf{I_{\text{shuffled}}}|$ do
16:     $(i,j)\leftarrow(I_{\text{shuffled}}[k],I_{\text{shuffled}}[l])$
17:      $(\mathbf{A^{*}_{ij}},\,\textit{outcome},\,r)\leftarrow\textsc{SolveA}^{*}\Bigl(i,\;j,\;\mathbf{G},\;\{a_{ig},a_{jg}\}_{g\in\mathbf{G}},\;B_{i},\;B_{j}\Bigr)$ $\triangleright$ Proposer $i$ finds best trade; $r$ is the responding agent (who accepts/rejects)
18:     if $\textit{outcome}=\textsc{accept}$ then $\triangleright$ Trade executed
19:      $\{a_{ig},\,a_{jg}\}_{g\in\mathbf{G}}\leftarrow\mathbf{A^{*}_{ij}}$
20:      $trades\_found\leftarrow\texttt{true}$
21:     end if
22:     $\textsc{BayesianUpdate}\bigl(i,\,j,\,r,\,\textit{outcome},\,B_{i},\,B_{j},\,\{B_{k}\}_{k\neq i,j}\bigr)$ $\triangleright$ All players update beliefs based on accept/reject
23:   end for
24:  end for
25:  $possible\_trades\leftarrow trades\_found$
26:end while
27:return $\mathbf{A}$
Explanation of key subroutines:
  • $\textsc{SolveA}^{*}(i,j,\dots)$:
    • Proposer’s optimization. Agent $i$ (the proposer) solves
      $\max_{(\mathbf{x},\mathbf{y})}\sum_{\mathbf{v}_{-i}}\mathbf{1}\Bigl\{\exists\,\text{receiver }r\neq i:\textsc{accept}\bigl(v_{r};\mathbf{x},\mathbf{y}\bigr)\Bigr\}\times\Bigl(u_{i}\bigl(h_{i}-\mathbf{x}+\mathbf{y}\bigr)-u_{i}\bigl(h_{i}\bigr)\Bigr)\times B_{i}(\mathbf{v}_{-i}),$
      where $\mathbf{x}$ are the chips given up by $i$ and $\mathbf{y}$ are the chips requested, $h_{i}$ denotes $i$’s current holdings, and $B_{i}$ is $i$’s belief over others’ valuations.
    • Receiver’s acceptance criterion. A potential responder $r$ accepts the trade $(\mathbf{x},\mathbf{y})$ myopically if
      $u_{r}\bigl(h_{r}-\mathbf{y}+\mathbf{x};\,v_{r}\bigr)\;>\;u_{r}\bigl(h_{r};\,v_{r}\bigr).$
      The subroutine returns $\mathbf{A^{*}_{ij}}$ (the updated allocation for $i$ and $j$), the outcome (accept or reject), and the identity of the actual responder $r$ (in case multiple receivers are considered, or if $j$ is always the designated responder, $r=j$).
  • $\textsc{BayesianUpdate}\bigl(i,j,r,\textit{outcome},B_{i},B_{j},\{B_{k}\}_{k\neq i,j}\bigr)$:
    • On acceptance: All agents discard any valuation states inconsistent with $r$ accepting $(\mathbf{x},\mathbf{y})$. Specifically,
      $v_{r}\text{ is retained only if }u_{r}(h_{r}-\mathbf{y}+\mathbf{x};v_{r})>u_{r}(h_{r};v_{r}).$
    • On rejection: All agents discard any valuation states for $r$ that would have accepted $(\mathbf{x},\mathbf{y})$, since that contradicts the observed rejection.
    • Then, each agent renormalizes its belief distributions $B_{k}$ to sum to 1 over the remaining (still-plausible) valuations.
This procedure repeats until no further beneficial trades can be found. The Bayesian updates ensure that, over time, agents refine their beliefs about each other’s valuations based on observed accept/reject decisions, thereby learning which trades are more likely to succeed.
Appendix B Human data collection
B.1 Implementation on Deliberate Lab
Platform.
The bargaining game was implemented via Deliberate Lab (Tsai et al. (2024)), an open-source platform for conducting group human-LLM social science experiments. We will publish the bargaining game implementation within the platform upon acceptance.
Participant interface.
Upon entering Deliberate Lab through a web link, participants proceed through a multi-stage experiment including informed consent, game instructions, comprehension checks, and payout information. Upon completing the final comprehension check, they wait in a “Lobby” stage for other participants. When three participants are in the lobby, they are sent an invitation to join a live bargaining game (Fig 6). Following the game, there is a post-game survey. For anonymity, we used a Deliberate Lab feature that assigns participants an anonymous animal avatar (e.g., “Bear”) as they join the experiment.
Figure 6: The game interface shown to participants. This shows a participant’s game state. The top left indicates that they are currently the Proposer, offering an interface to propose an offer. A table below displays all players’ chip volumes. To the right, the public ledger shows a history of game rounds, turns, and trade statuses.
Bargaining game user interface.
To submit trade proposals, participants used text and selection fields to specify number and type of chip, respectively. The interface automatically calculated and displayed projected total chip value if the proposal were to be accepted (Fig 6); it also prevented players from submitting invalid offers by disabling the "submit" button.
For accepting or rejecting proposals, the interface calculated and displayed the projected total chip value if the player were to accept the proposal. It also disabled acceptance for offers the player could not fulfill (Fig 7, bottom). All participants were shown a table of all players’ chip quantities and a log of actions (i.e., when trades were proposed, accepted, and rejected, and when each round and turn of the game changed). Turn order was determined randomly.
Figure 7: Accepting or rejecting a trade. The bottom image shows a disabled “accept” option.
Experimenter interface.
To conduct experiments, the experimenter used Deliberate Lab’s dashboard (Fig 8) to monitor participants as they progressed through the experiment’s information and comprehension check stages, then manually transferred participants into groups of three to play the bargaining game once the player was ready to be transferred. The experimenter was also able to send attention checks to participants (e.g., if several minutes elapsed without any visible progress) and remove them from the experiment (e.g., if attention checks went unanswered).
Figure 8: Experiment dashboard for managing participants. On the left, experimenters can view participants as they enter the experiment, send attention checks, and/or remove participants. On the right, experimenters can monitor the status of a selected participant and assign them to a cohort of three for gameplay.
Appendix C Post-game participant survey
C.1 Survey questions
Table 1: Post-game survey questions.
Survey question Response format
1 Please describe your strategy in the game in a few sentences. Freeform text
2 Please describe any experiences or background that may have influenced your performance in the game. Freeform text
3 On a scale from 1 to 10, how would you rate your trading strategy, where 1 is highly competitive (focused mainly on your own gains) and 10 is highly collaborative (focused on other players’s potential gains and mutual benefits? 1 through 10 scale; 1: "Not at all aggressive", 10: "Very aggressive"
4 On a scale from 1 to 10, how certain are you that you made the best possible trades during the experiment? 1 through 10 scale; 1: "Not at all confident", 10: "Very confident"
5 On a scale from 1 to 10, how satisfied are you with your final trading outcomes? 1 through 10 scale; 1: "Not at all satisfied", 10: "Very satisfied"
6 On a scale of 1 to 10, how would you rate the mental effort you put into today’s trading games, where 1 means ’I barely engaged in any thinking’ and 10 means ’I put in significant mental effort’? 1 through 10 scale; 1: "No mental exertion at all", 10: "Very high mental exertion"
7 Please provide any additional context on your answers above. Freeform text
8 Please help us to improve this experiment. How was your experience today? Were there any elements of the instructions or gameplay that you found confusing? Freeform text
C.2 Quantitative survey responses
Table 2: Summary statistics of self-assessment variables (mean and standard error). There was no significant change in certainty or exertion as game complexity increased.
Game variation Competitive Certainty Satisfaction Exertion
2 chip 5.57 (0.34) 5.57 (0.33) 6.68 (0.31) 7.11 (0.24)
3 chip 5.32 (0.34) 5.92 (0.33) 7.07 (0.30) 7.10 (0.25)
4 chip 5.75 (0.34) 6.04 (0.33) 7.46 (0.26) 7.46 (0.19)
Appendix D Additional clarification on strategic regret classifications
Figure 9: Regret behavior visualization: The top-left panel shows an overview trajectory in Fig 4. The top-right panel illustrates a “No Regret” scenario for P2 as a proposer. The bottom-left panel depicts “Forced Regret” for P3 as an acceptor. The bottom-right panel shows “Unforced Regret” for P2 as a decliner at Turn 8. Solid lines represent actual player paths, while dashed lines indicate counterfactual paths.
Here, we utilized counterfactual analysis to visually dissect the three kinds of regret behavior previously discussed in Section 6.2.2, as illustrated in Fig  9. In these visualizations, the player’s actual path of total chip value is depicted by a solid colored line, while the counterfactual path—representing the outcome had a different decision been made — is indicated by a gray dashed line.
Counterfactual definition:
The counterfactual path is constructed by simulating an alternative decision at a critical juncture (e.g., accepting a declined offer, declining an accepted offer, or making a different proposal) and projecting the subsequent value trajectory. This allows for a direct comparison between the actual outcome and what could have been.
No regret scenarios.
This scenario is demonstrated by Player 2 (P2) in the role of a proposer, as shown in the upper-right panel of Fig 9, focusing on Turn 9. Player 2 makes a proposal that is accepted, leading to an actual chip value of $15.3$. The counterfactual path, representing a scenario where P2 might not have made this proposal or made a less optimal one, results in a lower value of 14.9. In this “No Regret” instance, the player’s actual total value path is superior to or equal to the evaluated counterfactual path, indicating that the decision made was optimal given the available information and subsequent events.
Forced regret scenarios.
The lower-left panel of Fig 9 illustrates “Forced Regret” experienced by Player 3 (P3) as an acceptor from Turn 1 to Turn 2. P3 initially accepts an offer from P1 at Turn 1, resulting in an actual value of 10.8. However, a more lucrative offer from P2 becomes available in Turn 2. Due to the commitment made in Turn 1 (e.g., changed chip inventory), P3 is unable to capitalize on this subsequent, better opportunity. The counterfactual path shows that had P3 declined P1’s initial offer, they could have potentially accepted P2’s offer, leading to a higher value of 11.0.
Unforced regret scenarios.
The lower-right panel of  9 demonstrates Player 2 (P2) in the role of a decliner from Turn 8 to 9. P2 is presented with a profitable trade offer but chooses to decline it. P2’s actual chip value remains at 14.9 after this decision. The counterfactual path, however, demonstrates that had P2 accepted this profitable offer, their chip value would have increased to 15.0. This highlights an avoidable error: a superior option was available and actionable, but the agent failed to select it, leading to a sub-optimal outcome. Unlike forced regret, no prior decision prevented P2 from taking the better option; the regret stems directly from the choice made at that specific juncture.
Appendix E Surplus values and significance
Table 3: Scaled final surplus gain (mean and standard error). This is the ratio of observed total surplus to optimal surplus.
Game variation Human GPT-4o GPT-4o refined Gemini Gemini refined Bayesian agent
2 chip 0.60 (0.06) 0.69 (0.03) 0.68 (0.04) 0.42 (0.04) 0.45 (0.05) 0.74 (0.04)
3 chip 0.59 (0.04) 0.62 (0.03) 0.64 (0.03) 0.42 (0.03) 0.43 (0.03) 0.80 (0.03)
4 chip 0.54 (0.03) 0.54 (0.02) 0.58 (0.02) 0.37 (0.02) 0.33 (0.02) 0.73 (0.02)
Table 4: Two-sided t-test p-values comparing surplus values across populations (N=144 per game). Arrows indicate whether the column’s mean surplus was higher (↑) or lower (↓) than the row’s.
Human GPT-4o GPT-4o refined Gemini Gemini refined Bayesian agent
2 chip
Human 0.491 ↑ 0.597 ↑ 0.263 ↓ 0.066 ↓ 0.654 ↑
GPT-4o 0.302 ↓ 0.000 ↓ 0.000 ↓ 0.141 ↑
GPT-4o refined 0.000 ↓ 0.000 ↓ 0.414 ↑
Gemini 0.359 ↑ 0.000 ↑
Gemini refined 0.000 ↑
Bayesian agent
3 chip
Human 0.525 ↑ 0.021 ↑ 0.001 ↓ 0.000 ↓ 0.000 ↑
GPT-4o 0.164 ↑ 0.000 ↓ 0.004 ↓ 0.000 ↑
GPT-4o refined 0.000 ↓ 0.000 ↓ 0.001 ↑
Gemini 0.799 ↑ 0.000 ↑
Gemini refined 0.000 ↑
Bayesian agent
4 chip
Human 0.601 0.258 ↑ 0.000 ↓ 0.000 ↓ 0.000 ↑
GPT-4o 0.270 ↑ 0.000 ↓ 0.000 ↓ 0.000 ↑
GPT-4o refined 0.000 ↓ 0.000 ↓ 0.000 ↑
Gemini 0.003 ↓ 0.000 ↑
Gemini refined 0.000 ↑
Bayesian agent
Appendix F Aside on smaller model performance
We tested smaller models from both the Gemini and GPT families— Gemini-2.5-Flash and GPT-o4-mini—and found that their overall performance falls short of both human participants and larger models such as GPT-4o, particularly in terms of surplus generation. Even with prompt refinement, the average surplus gain remains substantially lower.
Procedurally, these small models also behave quite differently from their larger counterparts. As shown in Figure 10, proposals by Gemini-2.5-Flash and GPT-4-mini are heavily concentrated around a narrow region: nearly zero net surplus gain and a 1:1 trade ratio. Roughly 90% of offers consist of exchanging one chip for exactly one of another type (e.g., 1 red for 1 green). Loss-incurring proposals are virtually nonexistent—but so are high-surplus trades (e.g., net surplus >2>2). Prompt refinement slightly broadens the distribution but fails to shift the models away from this dominant 1:1 pattern.
This low-variance, risk-averse behavior is characteristic of the “lazy generation” effect reported in small language models (Lambert and Calandra, 2023), and reflects the narrow, template-driven action spaces observed in recent bargaining benchmarks for sub-10B parameter agents (Xia et al., 2024).
Figure 10: Comparison of the trade space (3-Chip) between small LLM agents (GPT-o4-mini and Gemini-2.5-flash).
F.1 Full strategic analysis visualization
Figure 11: This visualization follows the same structure as Figure 5, with additional Refined LLMs. Strategic differences between refined and out-of-the-box LLMs are empirically similar.
F.2 Supplementary trade space visualizations
Summary statistics of trade space values are in Table 5, with the 2-chip visualization in Figure 12 and 4-chip visualization in Figure 13. They exhibit similar behaviors as discussed in Section 6.2.1.
Figure 12: Comparison of the trade space (2-Chip) between Human participants, LLM agents (GPT-4o and Gemini-1.5 Pro), and Bayesian agents.
Figure 13: Trade space comparison (4-chip) between humans, LLM agents (GPT-4o and Gemini-1.5 Pro), and Bayesian agents.
Table 5: Summary statistics from trade space visualizations.
Game Player Acceptance Surplus Mean (SD) Ratio Mean (SD) Median [Surplus, Ratio]
2-chip Human Accepted 0.611 (1.031) 1.147 (0.652) [0.500, 1.000]
Rejected 1.088 (1.313) 0.735 (0.368) [0.800, 0.667]
GPT-4o Accepted 0.148 (0.617) 1.711 (0.976) [0.100, 1.500]
Rejected 0.414 (1.031) 0.715 (0.393) [0.100, 0.667]
Gemini-1.5 Pro Accepted -0.305 (0.913) 1.508 (0.721) [-0.100, 1.500]
Rejected 0.230 (1.317) 1.048 (0.475) [0.100, 1.000]
Bayesian Accepted 2.245 (1.252) 0.867 (0.315) [2.300, 0.900]
Rejected 2.058 (1.904) 0.636 (0.372) [1.500, 0.526]
3-chip Human Accepted 1.211 (1.317) 1.039 (0.387) [1.100, 1.000]
Rejected 1.830 (1.803) 0.840 (0.325) [1.300, 1.000]
GPT-4o Accepted 0.437 (0.838) 1.765 (0.945) [0.200, 1.600]
Rejected 0.345 (0.963) 0.918 (0.501) [0.200, 0.800]
Gemini-1.5 Pro Accepted 0.329 (0.899) 1.394 (0.610) [0.400, 1.500]
Rejected 0.528 (1.819) 1.075 (0.497) [0.600, 1.000]
Bayesian Accepted 3.173 (1.853) 0.944 (0.274) [3.200, 0.900]
Rejected 3.385 (2.725) 0.783 (0.355) [2.800, 0.700]
4-chip Human Accepted 1.331 (1.214) 1.173 (0.562) [1.200, 1.000]
Rejected 1.757 (1.438) 0.863 (0.288) [1.500, 1.000]
GPT-4o Accepted 0.369 (0.929) 1.745 (0.831) [0.200, 1.600]
Rejected 0.534 (0.981) 1.262 (0.518) [0.300, 1.250]
Gemini-1.5 Pro Accepted 0.403 (0.858) 1.438 (0.540) [0.400, 1.500]
Rejected 0.650 (1.187) 1.022 (0.377) [0.600, 1.000]
Bayesian Accepted 3.536 (1.793) 0.965 (0.251) [3.400, 0.900]
Rejected 4.251 (2.824) 0.738 (0.303) [3.500, 0.700]
Appendix G LLM simulation prompts
G.1 Rule explanation
How the game works:
The game consists of 3 rounds of trading. During each round, each player will have a turn to propose 1 trade. These turns are pre-determined in a random order and the order stays the same in each round.
Trade proposals:
To propose a trade, a player must:
1. Request a certain quantity of chips of a single color from any other player to get.
2. Specify a certain quantity of chips of a different color to give in return.
Trade rules:
Players cannot offer more chips than they currently hold. For example, if you only have 5 red chips, you cannot offer 6 red chips.
Players cannot trade chips of the same color. For example, you cannot trade red chips for red chips.
Trade completion:
When an offer is presented, all other active participants get a chance to accept or decline. Note: Active participants are those not currently making the offer.
Participants make their decisions simultaneously and privately. The participant who receives the offer is not dependent on who accepts the trade first. Some possible outcomes:
If no one accepts, the trade does not happen, and the turn ends.
If multiple participants accept, one accepting participant is chosen at random to complete the trade with the offering participant. This means that participants cannot choose who they trade with.
If only one participant accepts, the trade will happen.
Key points to remember:
In each round, each player gets to propose one trade and respond to other players trades
You can only propose trades between different colored chips, and cannot offer to give a chip amount that you do not have
When multiple players accept a trade, the trading partner is randomly selected
G.2 Proposer’s prompt
We used a Chain-of-Thought approach to elicit LLM agent’s trade proposals:
You are {{name}}.
Your valuations of the different types of chips are: {{preference_description}}.
You now have the following amounts of each chip: {{item}}.
The conversation history so far is {{history}}.
REMEMBER, to propose a trade, you must:
Request a certain quantity of chips of a single color from any other player to get
Specify a certain quantity of chips of a different color to give in return
REMEMBER you have the following amounts of each chip: {{item}}.
Your goal is to make as much money as possible. The trades, you choose to make to accomplish this, are up to you.
As a part of making money you must be rational - do not propose a trade in which you lose money. The value of a trade to you is the difference between the total value of chips you receive (quantity x your valuation) minus the total value of chips you give up (quantity x your valuation). Only propose trades that give you positive value.
In short, your trades should be both incentive compatible and incentive rational.
Your response must use these EXACT tags below. The response should include nothing else besides the tags, your trade offer, and your reasoning. The text between tags should be concise.
<REASONING>
[Provide your concise reasoning in a few sentences, e.g. To gain more surplus, I want more xxx chips]
</REASONING>
<CHECK>
[check if you have sufficient chips to trade. If you have n green chips, you can at most give n green chips. If you dont want to trade, you can ask for a large amount of chips that no one can afford]
[…]
<GET_COLOR> Color, e.g. red</GET_COLOR>
<GET_QUANTITY> quantity, e.g. n </GET_QUANTITY>
<GIVE_COLOR> Color, e.g. red</GIVE_COLOR>
<GIVE_QUANTITY> quantity, e.g. n </GIVE_QUANTITY>
G.3 Receiver’s prompt
You are {{name}}.
Your valuations of the different types of chips are: {{preference_description}}.
You now have the following amounts of each chip: {{item}}.
The conversation history so far is {{history}}.
You have an offer. {{proposer}} is offering to give {{give}} and get {{get}} in return.
If you make this trade, your total wealth will change by: {{delta_surplus}}
Now, you need to decide whether to accept or decline.
Your response must use these EXACT tags below. The response should include nothing else besides the tags, your choice to accept or decline, and your reasoning. The text between tags should be concise.
<REASONING>
[Provide your concise reasoning in a few sentences.]
</REASONING>
<CHOICE>Yes or No </CHOICE>
G.4 Refined Proposing Prompt
The refined prompt first generates multiple trades proposals and uses another LLM query to choose among them.
G.4.1 Multiple proposal generation prompt
You are {{name}}.
Your valuations of the different types of chips are: {{preference_description}}.
You now have the following amounts of each chip: {{item}}.
The conversation history so far is {{history}}.
REMEMBER, to propose a trade, you must:
Request a certain quantity of chips of a single color from any other player to get
Specify a certain quantity of chips of a different color to give in return
REMEMBER you have the following amounts of each chip: {{item}}.
Your goal is to make as much money as possible. The trades, you choose to make to accomplish this, are up to you.
As a part of making money you must be rational - do not propose a trade in which you lose money. The value of a trade to you is the difference between the total value of chips you receive (quantity $\times$$ your valuation) minus the total value of chips you give up (quantity $\times$$ your valuation). Only propose trades that give you positive value.
In short, your trades should be both incentive compatible and incentive rational.
You can trade as many chips as you want in a single turn, assuming you have that many. Do not feel constrained to only trade a single chip at a time.
Propose 3 different good trade ideas, so a next step can decide on the best trade of the ones you propose here.
Your response must use these EXACT tags below. The response should include nothing else besides the tags, your trade offer, and your reasoning. The text between tags should be concise.
Repeat the below tags once for each of the trade ideas you propose.
<REASONING>
[Provide your concise reasoning in a few sentences, e.g. To gain more surplus, I want more xxx chips]
</REASONING>
<CHECK>
[check if you have sufficient chips to trade. If you have n green chips, you can at most give n green chips. If you dont want to trade, you can ask for a large amount of chips that no one can afford]
[…]
<GET_COLOR> Color, e.g. red</GET_COLOR>
<GET_QUANTITY> quantity, e.g. n </GET_QUANTITY>
<GIVE_COLOR> Color, e.g. red</GIVE_COLOR>
<GIVE_QUANTITY> quantity, e.g. n </GIVE_QUANTITY>
G.5 Choosing from multiple possible proposals prompt
You are {{name}}.
Your valuations of the different types of chips are: {{preference_description}}.
You now have the following amounts of each chip: {{item}}.
The conversation history so far is {{history}}.
REMEMBER, to propose a trade, you must:
Request a certain quantity of chips of a single color from any other player to get
Specify a certain quantity of chips of a different color to give in return
REMEMBER you have the following amounts of each chip: {{item}}.
Your goal is to make as much money as possible. The trades, you choose to make to accomplish this, are up to you.
As a part of making money you must be rational - do not propose a trade in which you lose money. The value of a trade to you is the difference between the total value of chips you receive (quantity $\times$$ your valuation) minus the total value of chips you give up (quantity $\times$$ your valuation). Only propose trades that give you positive value.
In short, your trades should be both incentive compatible and incentive rational.
You can trade as many chips as you want in a single turn, assuming you have that many. Do not feel constrained to only trade a single chip at a time.
Below are three trade ideas you have proposed. Please pick the best trade proposal from the ones below that you will propose to the group.
Do not change anything about the trade you have selected from the ideas below.
Your response must use these EXACT tags below. The response should include nothing else besides the tags and content of your selected trade offer including its reasoning.
<REASONING>
[Provide your concise reasoning in a few sentences, e.g. To gain more surplus, I want more xxx chips]
</REASONING>
<CHECK>
[check if you have sufficient chips to trade. If you have n green chips, you can at most give n green chips. If you dont want to trade, you can ask for a large amount of chips that no one can afford]
[…]
<GET_COLOR> Color, e.g. red</GET_COLOR>
<GET_QUANTITY> quantity, e.g. n </GET_QUANTITY>
<GIVE_COLOR> Color, e.g. red</GIVE_COLOR>
<GIVE_QUANTITY> quantity, e.g. n </GIVE_QUANTITY>
Proposed trade ideas to choose from:
{{proposed}}