### v4

# Strategic Tradeoffs Between Humans and AI in Multi-Agent Bargaining

Abstract.
Markets increasingly accommodate large language models (LLMs) as autonomous decision-making agents. As this transition occurs, it becomes critical to evaluate how these agents behave relative to their human and task-specific statistical predecessors. In this work, we present results from an empirical study comparing humans (N=216), multiple frontier LLMs, and customized Bayesian agents in dynamic multi-player bargaining games under identical conditions. Bayesian agents extract the highest surplus with aggressive trade proposals that are frequently rejected. Humans and LLMs achieve comparable aggregate surplus within their groups, but exhibit different trading strategies. LLMs favor conservative, concessionary proposals that are usually accepted by other LLMs, while humans propose trades that are consistent with fairness norms but are more likely to be rejected. These findings highlight that performance parity—a common benchmark in agent evaluation—can mask substantive procedural differences in how LLMs behave in complex multi-agent interactions.
multi-agent bargaining; negotiation; large language models; autonomous agents; strategic behavior; experimental study; Bayesian agents; fairness norms
journalyear: 2026copyright: ccconference: 31st International Conference on Intelligent User Interfaces; March 23–26, 2026; Paphos, Cyprusbooktitle: 31st International Conference on Intelligent User Interfaces (IUI ’26), March 23–26, 2026, Paphos, Cyprusdoi: 10.1145/3742413.3789078isbn: 979-8-4007-1984-4/2026/03ccs: Human-centered computing Empirical studies in HCIccs: Computing methodologies Multi-agent systemsccs: Theory of computation Algorithmic game theory and mechanism designccs: Computing methodologies Artificial intelligence
1. Introduction
For actions with real social or economic consequences, humans have traditionally been the decision-makers, while statistical models have handled narrowly-scoped, well-defined tasks. As large language models (LLMs) have advanced, they have begun to upend this paradigm (Eloundou et al., 2024; Shahidi et al., 2025), exhibiting strong zero- and few-shot performance across many domains (Brown et al., 2020). For example, call-center inquiries were once handled solely by human agents, then fronted by scripts and decision trees, and are now increasingly triaged by LLMs before reaching humans (Brynjolfsson et al., 2025). Similar shifts are emerging across other interpersonal tasks. LLMs are used to negotiate contracts (Van Hoek et al., 2022), coordinate trade (Corvin, 2024), and facilitate e-commerce (PYMNTS, 2025; Gaarlandt et al., 2025). This proliferation continues even though LLMs can fail to generalize in unexpected ways (Vafa et al., 2024). As they take on higher-stakes decisions, understanding their behavior relative to humans and previous generations of statistical models becomes essential—not only in terms of outcomes, but also in terms of alignment with procedural strategies and norms.
Evaluating these tradeoffs requires assessing different types of agents in comparable yet sufficiently complex settings. This is difficult in practice. Across real-world applications, populations, incentives, and interfaces vary, making it hard to attribute observed differences to the agent rather than the setting. Static QA-style benchmarks—still the dominant LLM evaluation paradigm (Hendrycks et al., 2021)—fail to capture the dynamic, multi-agent settings that characterize many real-world interactions. Indeed, recent work has called for more realistic evaluation environments that reflect the social and strategic complexities of the real-world (Raman et al., 2024; Goktas et al., 2025).
We address this challenge by introducing an incentivized, multi-player bargaining game. The game is designed to enable controlled comparisons among LLMs, humans, and bespoke statistical agents under identical conditions. In our setting, players are endowed with colored chips associated with fixed, private valuations, and they negotiate trades to maximize individual surplus, defined as the value of their final chip bundle minus the value of their initial endowment. Crucially, preferences are fully induced (Smith, 1976): endowments and valuations are randomly drawn from a known distribution. This design ensures that differences in outcomes or behavior are attributable to differences in the agents themselves, rather than differences in underlying preferences.
This game has several features that make it well-suited for comparing agent types. First, its specific formulation is novel, providing a meaningful out-of-distribution domain for LLM agents. Second, despite the controlled structure of the game, bargaining under partial information admits no Pareto-efficient, incentive-compatible solution (Myerson and Satterthwaite, 1983), forcing agents to navigate genuine tradeoffs between self-interest and cooperation. Third, we can vary strategic complexity by adjusting the number of chip types in players’ endowments. Finally, while the game is not itself incentive-compatible, it does admit a computable Pareto-optimal upper bound, which can benchmark performance.
We conduct an empirical study of how different agent types play this game under identical conditions. First, we run an incentivized online study with human participants (N=216N=216). [^1] Humans are paid in USD per their gains from trade beyond their initial endowment. We then replicate the same randomized initial conditions in simulation with (i) LLM-based agents (GPT-4o and Gemini 1.5 Pro) and (ii) specialized Bayesian agents, enabling controlled comparisons across the three agent types. This setup enables an examination of two central research questions:
  1. RQ1:
    Performance: How does the performance of LLM agents compare to bespoke statistical agents and humans?
  2. RQ2:
    Process: How do the trading strategies and resulting group dynamics of LLM agents compare to those of other agents?
Figure 1. Overview of the bargaining game. Left: Game setup showing chip endowments and valuation structure. Center: Gameplay mechanics for proposing and accepting offers. Right: The three agent types evaluated in this study.
Three-panel overview of a three-player bargaining game. Left: players receive chips of 2–4 colors; chip counts are public; green has a shared value, and other colors have private random values. Center: players propose in a fixed cycle; the two others accept or reject, yielding no deal, a bilateral deal, or a random tie-break if both accept. Right: agents include humans, LLMs (Gemini 1.5 Pro, GPT-4o), and heuristic baselines (myopic, Bayesian).
We find that the game proves nontrivial: across all games and complexity levels, no agent type—human or artificial—achieves more than 80% of the Pareto bound on average. The agents’ performance varied substantially, and the processes by which they traded were remarkably different. Bayesian agents achieved the highest average surplus (up to 80% of the Pareto bound) by proposing highly self-interested trades that were often rejected. LLM agents were capable of achieving similar surplus (62%) to their human counterparts (59%). [^2] These statistics are for the medium-complexity game variation (3-chip). See Section 3.
Despite comparable outcomes, strategic behaviors between humans and GPT-4o agents diverged substantially. Humans offered trades that appeared more “fair,” while LLMs proposed more conservative trades that gave other LLMs within their group substantial surplus and were accepted at higher rates. Taken together, these results highlight the limitations of outcome-driven evaluation, as the similar surplus gains by humans and LLMs masked consequential differences in strategic behavior and procedural alignment.
In summary, this work makes three core contributions:
  1. 1.
    A dynamic multi-agent environment for comparative evaluation. We design and implement an open-source, incentivized bargaining game that enables ceteris paribus comparisons of behavior across agent types, with game variations of increasing complexity. The game is extensible to agent types not studied here (e.g., other LLMs). We also provide a linear programming formulation of the Pareto-optimal surplus as a theoretical benchmark for performance.
  2. 2.
    An empirical sample of human, LLM, and Bayesian trading strategies. We characterize how each agent type negotiates within a bargaining scenario under identical conditions: Bayesian agents aggressively maximize surplus and frequently reject offers; humans make strategic concessions motivated by fairness norms; LLMs negotiate conservatively and accept most offers.
  3. 3.
    Evidence that outcome similarity can mask meaningful differences in behavior. We show that aggregate outcomes (e.g., total surplus) can conceal substantial differences in strategies, interaction patterns, and normative alignment. This finding highlights the need to move beyond performance-based metrics for the responsible deployment of LLMs in high-stakes real-world settings.
2. Related Work
Recent advances in LLMs’ capabilities have dramatically broadened the scope of tasks that they can perform in scenarios involving collective reasoning, including understanding of social context, preference elicitation, strategic pricing, and natural-language negotiation (Fish et al., 2024; Soumalias et al., 2025; Tessler et al., 2024; Vaccaro et al., 2025; Lubars and Tan, 2019; Bakhtin et al., 2022; Kramár et al., 2022; Deng et al., 2024; Allouah et al., 2025). LLMs have also demonstrated the capacity to perform economic tasks, replicating and predicting behaviors in bargaining, auctions, and other games (Aher et al., 2022; Argyle et al., 2022; Horton, 2023; Manning et al., 2024; Manning and Horton, 2025; Xie et al., 2024; Zhu et al., 2024; Park et al., 2023; Qian et al., 2025a).
Yet demonstrating that LLMs can perform economic tasks leaves open the question of how they do so. Imas et al. (2025) study an experimental marketplace in which humans write prompts for AI agents that negotiate on their behalf. They find that, contrary to predictions of homogenized outcomes, agentic interactions exhibit greater dispersion than human-to-human negotiations. This suggests that prompts are not neutral instructions but carriers of human heterogeneity. Our work complements theirs: while they focus on how human characteristics propagate through delegation, we focus on comparing how different agent types behave when facing identical induced preferences, isolating strategic differences rather than principal heterogeneity.
These findings underscore a broader point: AI agents warrant behavioral study in their own right. Rahwan et al. (2019) argue that the behavioral study of AI agents necessitates its own scientific research agenda. This is particularly critical in mixed-motive and multi-agent settings, where alignment is not static but dynamically negotiated. Recent work on virtual bargaining suggests that effective coordination requires agents to mentally simulate implicit social contracts rather than solely maximizing reward functions (Chater, 2023). Furthermore, Bansal et al. (2019) highlight a “performance-compatibility tradeoff”: updates that increase an agent’s theoretical accuracy can paradoxically degrade team performance if they violate the user’s mental model of the agent’s error boundaries.
More generally, a broad theme in recent findings is that appropriate deployment depends not only on system capabilities (Jahani et al., 2025), but also on alignment with user expectations and social norms (Mozannar et al., 2023; Agarwal et al., 2023; Mullainathan and Obermeyer, 2021; Palminteri et al., 2024; Kapania et al., 2022; Qian and Wexler, 2024). Addressing these alignment gaps requires evaluating agents in dynamic contexts with social consequences—a need recognized by recent calls for rigorous methods to steer and assess agent behavior (Raman et al., 2024; Goktas et al., 2025). We contribute to this agenda by studying how humans, LLM-based agents, and Bayesian models behave in a dynamic bargaining task, offering a comparative lens on alignment in strategic social interactions.
3. The Bargaining Game
Figure 1 provides an overview of the game. Below, we describe the setup, information structure, and gameplay.
Setup.
A game consists of 3 players (any $n>2$ suffices) and $k$ distinct types of colored chips, where $k\in\{2,3,4\}$ varies across experimental conditions. [^3] Game complexity is increasing in $k$. Proof provided in Appendix A.1. Players begin with 10 chips of each type. One chip type is always green, which serves as a shared numeraire publicly valued at \$0.50 by all players. The remaining chip types (drawn from red, blue, and purple, depending on $k$) have private valuations: each player independently draws a value for each non-green chip type uniformly from [\$0.10, \$1.00]. While players do not know the valuations other players hold for non-green chips, they do know the distribution from which these values are drawn. Crucially, this setup provides players with an incentive to trade because the possible mutual gains may be substantial. For example, if one player values blue chips at \$0.20 and another values them at \$0.90, trading blue chips for green chips (the shared numeraire) allows both players to capture part of the \$0.70 valuation gap.
Gameplay.
Players take turns in a fixed and initially randomized order, cycling through until each player has proposed three times (9 total proposals for 3 players). On their turn, a player proposes a trade: offering chip(s) of one color in exchange for chip(s) of another. The non-proposing players simultaneously decide whether to accept or reject. If both accept, one is chosen at random to complete the trade. Players may only propose or accept trades that they can fulfill with their current inventory. At the end of the game, each player’s payout equals the difference in value between their final and initial chip holdings, floored at zero. Throughout the game, players can see the number of chips each player holds, although not their opponents’ valuations, along with a live history of all previous proposals and trades.
4. A Pareto-Optimal Upper Bound
Next, we provide a theoretical upper bound analysis to benchmark agent performance in our empirical results. Due to the complexity of identifying a tractable, exact equilibrium in this setting, we instead compute a Pareto-optimal allocation via linear programming.
4.1. Notation
Consider a game M = (I, G, v, a) with a set of agents I and chips G. Each agent $i$ has a personal valuation $v_{ig}$ for each good $g$. These valuations are static, independent across agents, and complete for each agent-good pair. All agents have an initial allocation $a^{0}_{ig}$ of each chip, which indicates the amount of chip $g$ allocated to agent $i$. Given these allocations, an agent’s welfare $w_{i}$ is the sum of their valuations weighted by their allocations: $w_{i}=\sum_{g\in G}v_{ig}a_{ig}$. The total welfare w is thus the sum of all individual welfare: $\textbf{w}=\sum_{i\in I}w_{i}=\sum_{i\in I}\sum_{g\in G}v_{ig}a_{ig}$. An individual’s surplus (and corresponding final payout) is defined as the difference between their final and initial welfare.
4.2. Efficient Allocations
In our game, agents trade chips to improve their surplus. Given their initial endowments $a^{0}_{ig}$, mutually beneficial trades may exist that increase total welfare. We provide an upper bound to the obtainable total welfare, assuming the Pareto condition that no player ends with worse utility than they began with.
Definition 1.
An optimal allocation of chips is any set of allocations $\mathbf{A^{*}}=\{a^{*}_{ig}\ |\ i\in I,g\in G\}$ that solves:
$\operatorname*{\arg\!\max}_{\mathbf{A}}\sum_{i\in I}\sum_{g\in G}v_{ig}a_{ig}$
subject to:
  1. (i)
    Conservation of goods: $\sum_{i\in I}a_{ig}=\sum_{i\in I}a^{0}_{ig},\forall\ g\in G$
  2. (ii)
    Pareto improvement: $\sum_{g\in G}v_{ig}a_{ig}\ \geq\ \sum_{g\in G}v_{ig}a^{0}_{ig},\forall\ i\in I$.
  3. (iii)
    Non-negativity: $a_{ig}\geq 0\forall\ i\in I,g\in G$.
Under a centralized setting with full information, $\mathbf{A^{*}}$ represents a Pareto-efficient allocation that maximizes aggregate utility without making any agent worse off. While this optimal allocation $\mathbf{A^{*}}$ may not be unique, the maximum total welfare $\mathbf{w^{*}}$ is unique. This optimization problem can be solved efficiently using interior point methods. Its computational complexity is bounded by $O\bigl((|I||G|)^{3.5}\bigr)$. Note that this upper bound is unlikely to be achieved in practice, as no mechanism for decentralized bargaining with private information can guarantee both Pareto efficiency and incentive compatibility (Myerson and Satterthwaite, 1983). [^4] See Appendix A.2 for a proof that players have no dominant strategy in our game.
5. Experimental Setup
We implemented the game using Deliberate Lab (Qian et al., 2025b), an open-source platform for conducting multi-agent experiments. Our setup ensures full reproducibility, allowing for exact replication of game mechanics for benchmarking of future LLMs. Figures 6 and  B.1 provide visualizations of the game interface on Deliberate Lab.
In our experimental design, human participants were randomly assigned to groups of three to complete either the 2-chip, 3-chip, or 4-chip game. Humans also completed an end-of-task survey that asked them qualitative questions about their experience playing the game. [^5] Full survey questions are provided in Appendix C.
These human game instances—defined by the unique initial chip endowments within a group—were then replicated in isolation using LLM and Bayesian agents. To facilitate a direct performance comparison, all groups remained homogeneous; no interactions occurred between different agent types (e.g., humans played exclusively against humans, and LLMs against LLMs).
5.1. Human Subjects
We recruited 216 U.S.-based participants from Prolific to complete the bargaining game on Deliberate Lab. [^6] Institutional IRB approval and informed consent were obtained; we applied no inclusion criteria for recruitment. Each participant completed two games over a 30-minute session. To incentivize play, one game was randomly selected for payment, and participants received their surplus—the difference between their final and initial chip values—as a bonus in addition to a \$4 base payment. For example, a player whose chips increase in value from \$10 to \$14 earns a \$4 surplus bonus. The average total payout was \$12.24 $\pm$ \$6.12. Each participant completed two games with different chip valuations and profiles. For each of the 2-, 3-, and 4-chip game variants, 24 groups of three players played 2 games each with 9 turns, yielding 1,296 trade proposals across 144 games. Participants were not informed whether they were matched with the same partners across games. The two games are evaluated independently, as ordering and learning effects proved minimal. [^7] Learning effects are discussed in Appendix F.1.
5.2. LLM Agents
We replicated each human game with two types of LLM agents. First, we wanted to evaluate how LLMs perform “out of the box,” constructing a baseline agent using intentionally simple prompts adapted directly from the human instructions. Second, we wanted to see if we could improve performance by prompting the LLM to reason more deeply about the strategic context. To do this, we implemented a refined agent that first generates multiple candidate proposals and then uses a second LLM call to select among them—mimicking the slower, more deliberate “Type 2” decision-making process described by Kahneman (2011) and applied to LLMs by Furniturewala et al. (2024). We tested both prompting strategies on OpenAI’s gpt-4o-2024-05-13 (gpt-4o, hereinafter) (OpenAI et al., 2024) and Google’s gemini-1.5-pro-001 released on 2024-05-23 (Gemini 1.5 Pro) (Team et al., 2024), which were the latest full-capacity models from each family at the time of writing. [^8] We also tested (Gemini 2.5 Flash, GPT-o4-mini), but they could not reliably propose valid trades. See Appendix G for further details. Model temperature was set to 0.5.999A lower temperature risked agents repeatedly sampling a narrow subset of trades (Lambert and Calandra, 2023). A higher temperature risked invalid trades (Arora et al., 2024). Prompts were designed using chain-of-thought (Wei et al., 2022), and simulations were implemented using the open-source EDSL library (Horton et al., 2024).
Importantly, the replications were identical to the human games in terms of initial conditions and gameplay. At each turn, an LLM agent received the same information as its human counterpart, including private chip valuations, public chip holdings, and trade history. One agent proposed a trade per round, and the others evaluated it.
Figure 2. Surplus trajectories across game complexities by agent type. Each blue line traces a game’s surplus as a fraction of the Pareto-optimal benchmark; black lines show average trajectories with 95% confidence bands. The dashed line marks the theoretical ceiling (ratio = 1), and each panel displays the final average at Turn 9. Final surplus values are reported in Table 3.
Surplus trajectories across game complexities by agent type. Each blue line traces a game’s surplus as a fraction of the Pareto-optimal benchmark; black lines show average trajectories with 95% confidence bands. The dashed line marks the theoretical ceiling (ratio = 1), and each panel displays the final average at Turn 9.
5.3. Bayesian Agents
We include a Bayesian agent as a task-specific benchmark—an explicitly programmed agent that optimizes within the known structure of the game. We summarize the algorithm here and provide full details in Appendix A.3. Each agent $i$ forms a joint belief about the valuations of all other players, $\mathbf{v}_{-i}$. When agent $i$ proposes a trade $(x_{g},y_{r})$ – offering $x$ chips of type $g$ in exchange for $y$ chips of type $r$ – they choose a trade to maximize their expected payoff:
Figure 3. Trading patterns in the 3-chip game. Top: Net surplus change for the proposer. Values to the right of the dashed line represent positive surplus to the proposer. Bottom: Chip trading ratio (number of chips offered to requested). Accepted trades are shown in green, trades that were not accepted (for any reason) are in red. Values near the dashed line mark balanced exchange (a 1:1 ratio of chips).
Trading patterns in the 3-chip game. Top: Net surplus change for the proposer. Values to the right of the dashed line represent positive surplus to the proposer. Bottom: Chip trading ratio (number of chips offered to requested). Accepted trades are shown in green, trades that were not accepted (for any reason) are in red. Values near the dashed line mark balanced exchange (a 1:1 ratio of chips).
$\max_{(\mathbf{x},\mathbf{y})}\sum_{\mathbf{v}_{-i}}\mathbf{1}\Bigl\{\exists\,j\neq i:\text{accept}\bigl(v_{j},x_{g},y_{r}\bigr)\Bigr\}\;\times\;\Delta u_{i}(v_{i},x_{g},y_{r})\;\times\;B_{i}\bigl(\mathbf{v}_{-i}\bigr),$
where $\text{accept}\bigl(v_{j},x_{g},y_{r}\bigr)$ indicates that agent $j$ with valuation $v_{j}$ would accept the trade, $\Delta u_{i}(v_{i},x_{g},y_{r})=v_{i,r}y_{r}-v_{i,g}x_{g}$ is agent $i$’s change in utility and $B_{i}(\mathbf{v}_{-i})$ is player $i$’s belief (probability) that the other agents’ valuations are $\mathbf{v}_{-i}$. This agent assumes that their opponent’s acceptance decision is myopically rational, i.e., a receiver with valuation $v_{j}$ accepts if and only if they have sufficient chips for the trade and their expected utility is positive, i.e., $v_{j,g}x_{g}-v_{j,r}y_{r}>0$.
All agents update their priors after observing acceptance decisions; they discard valuation states $\{v_{j}\}$ that contradict the observed event, assuming myopically rational opponents, and re-normalize the remaining probabilities. When the trade is accepted, the proposer keeps only valuations that would be accepted in its distribution over the chosen receiver. The receiver keeps only valuations that would be accepted in its distribution over the proposer. Bystanders update both priors accordingly. When there is no trade, all players discard valuations that conflict with the observed rejections.
6. Results
We organize our findings around the two research questions. For RQ1 (performance), we compare aggregate surplus across agent types. For RQ2 (process), we examine trading patterns both at the turn- and game-level. We additionally report qualitative feedback from human participants on their strategies and experiences.
6.1. Performance
Figure 2 shows how total group surplus—the sum of all players’ gains from trade—evolves over each game’s nine turns. Bayesian agents consistently achieved the highest surplus across all game variants, reaching 74–80% of the optimum. GPT-4o performed comparably to human participants, while Gemini 1.5 Pro achieved lower aggregate surplus (42–45% vs. 54–69% for humans and GPT-4o). The difference in performance between GPT-4o and the humans is negligible. The refined prompting variants did not substantially improve performance for either LLM family.
Beyond the final surplus, several patterns stand out in these trajectories. Bayesian agents show notably concave paths—rapid early gains that flatten as obvious trades are exhausted. In contrast, human and LLM trajectories rise more linearly, with Gemini 1.5 Pro showing the slowest average gains. As game complexity increased from two to four chip types, all agent types captured a smaller share of the optimal surplus, though Bayesian agents remained the most robust to this increase.
6.2. Procedural Alignment
Given that humans and GPT-4o agents achieved comparable surplus, we now ask whether they arrived there through similar processes. In this section, we evaluate strategic differences between the human, GPT-4o, and Bayesian agents. We analyze behavior at two levels, since agents may accept short-term losses to position themselves for later gains, or conversely, capture early value at the cost of foreclosing better opportunities. We first examine the structure of individual proposals at the turn level—what each trade would yield if accepted, without considering future opportunities. We then evaluate decisions retrospectively—whether they were optimal given the opportunities that arose later in the game.
6.2.1. Turn-level analysis
To characterize trading behavior at the turn level, we examine two features of each proposal: (i) the net surplus it would generate for the proposer, and (ii) the ratio of chips offered to chips requested.
Figure 3 shows these distributions for the 3-chip game.1010102- and 4-chip results are qualitatively similar; figures provided in Appendix F.2. Figure values are provided in Table 5. The top row shows that humans generally proposed trades that would generate positive surplus, though some yield net-zero or negative returns. The bottom row shows that human proposals tend to be fair in the number of chips exchanged. Both LLM families tend to propose trades that generate little surplus and more often result in a personal loss. On the chip trading ratio plot, a noticeable tail extends above the 1:1 line, indicating a tendency to offer more chips than requested (ratios up to 5:1).
By construction, Bayesian agents propose trades that give them surplus greater than zero—all density lies right of zero in the surplus plot. [^11] Density mass below zero surplus for the Bayesian agents is an artifact of Gaussian smoothing. None of the Bayesian proposals generate a negative surplus by construction. In the chip trading ratio row, Bayesian proposals cluster below 1:1 (ratio < 1); that is, they request more chips than they offer. Such proposals are frequently rejected.
These trading patterns reveal clear strategic differences. Humans offer trades suggesting a consideration for balanced, fair exchanges (Danz et al., 2022). LLMs propose highly acceptable trades consistent with a “concessionary posture”—offering many chips in return for a few. Bayesian agents adopt a value-extractive strategy (Fisher et al., 2011), asking to receive more than they give, at the cost of frequent rejections. To evaluate whether any of these strategies were optimal ex-post, we next apply a regret-minimization framework.
Figure 4. Example of player chip value trajectories over nine turns of a trading game (3-chip). To avoid speculation, we do not analyze unobserved strategies. For example, we do not analyze Player 1’s proposal on Turn 4 (an unaccepted trade) or Player 3’s acceptance on Turn 5 (accepting a negative surplus offer).
Example of player chip value trajectories over nine turns of a trading game (3-chip). To avoid speculation, we do not analyze unobserved strategies. For example, we do not analyze Player 1’s proposal on Turn 4 (an unaccepted trade) or Player 3’s acceptance on Turn 5 (accepting a negative surplus offer).
6.2.2. Retrospective analysis
We now evaluate decisions retrospectively, asking whether they were optimal given the opportunities that actually materialized later in the game. At each turn, a player takes one of three actions: proposing, accepting, or declining a trade. We classify each action by its regret properties:
  • No regret: There was no alternative move in a later round that would have yielded higher surplus.
  • Forced regret: A higher-surplus option arose later, but earlier decisions (such as premature trades or overcommitting inventory) prevented the agent from acting on it.
  • Unforced regret: A higher-surplus move was available to take (i.e., no budget or chip count constraints), yet the agent failed to take it.
We analyze how these regret classifications apply to the actions of the different agent types. To avoid speculative modeling of unobserved strategies, we restrict our analysis to a subset of trades where the agent’s intent is clear. First, we only consider myopically rational transactions for proposers, acceptors, or decliners. [^12] That is, only transactions yielding positive surplus. Note that Bayesian agents satisfy this by design. Interpreting irrational trades would require assumptions we cannot credibly test, as they may be driven by factors such as reciprocity, signaling, or cognitive error. Second, when evaluating proposer behavior, we restrict our regret analysis to focus on accepted trades. The act of rejecting an offer could be a mistake, but it could also be strategic signaling. While acceptances can also stem from participant error, they provide observable realized gains that allow for objective regret calculation. Third, “acceptors” include those who intended to accept a trade, not only those who were selected to trade. Recall that this occurs when both other players in a game accept the trade; one is randomly chosen.
Figure 4 shows an example of a player’s valuations as they trade over the course of a single game. On the left, private valuations are shown: Player 3 values blue chips at \$0.30 and red chips at \$0.10. On turn 1, Player 1 proposes to give 9 blue chips to receive 9 red chips in return. Player 3 is an acceptor and ends the turn with 19 blue chips and 1 red chip, netting a \$1.80 surplus. However, in turn 2, Player 2 offers a more lucrative offer: 10 blue chips for 10 red chips, which would give Player 3 a higher surplus of \$2.00 if they could accept it. However, Player 3 does not have enough red chips to clear the Turn 2 trade. Therefore, we characterize their action in Turn 1 as forced regret by the acceptor.
On turn 7, Player 2 makes an unforced regret decline. Player 1 offers to give two green chips for three blue chips. Player 2 has enough chips to clear this trade at this point; if Player 2 were to accept, they would gain $0.5*2-0.3*3=\mathdollar 0.10$. However, they decline the profitable trade and do not recover the profits later in the game. [^13] We do not classify proposals as unforced regret; agents may send intentionally conservative offers to elicit signals on market dynamics.
On turn 8, Player 2 makes a no regret proposal. Player 2 offers to give six red chips to receive eight green chips, resulting in a surplus of $0.5*8-0.6*6=\mathdollar 0.40$. This was the best deal they could have made, as there were no better trade opportunities in the future. Additional details on the regret taxonomy are provided in Appendix D.
Figure 5 counts these actions across agent types and game complexity. Each panel corresponds to a game variant (2, 3, or 4 chips); rows group agents by role (proposer, acceptor, decliner). Green bars indicate no-regret actions, blue bars unforced regret, and red bars forced regret. Because we restrict our analysis to unambiguous actions, these bars reflect the frequency of classifications within the filtered sample. The relative proportions of color within each bar, rather than the absolute bar length, provide the relevant comparison of strategic behavior across agent types.
Figure 5. Counts of actions by regret category across agent types and game complexity. Green is no regret, blue is unforced regret, and red is forced regret. Rows group agents by role (proposer, acceptor, decliner). Note: declining cannot cause forced regret since it does not commit chips.
Counts of actions by regret category across agent types and game complexity. Green is no regret, blue is unforced regret, and red is forced regret. Rows group agents by role (proposer, acceptor, decliner). Note: declining cannot cause forced regret since it does not commit chips.
Several patterns emerge. The proportion of green (no regret) in Bayesian proposals relative to other classifications is high. That is, they consistently identify optimal trades. In contrast, human and LLM proposers show substantial relative unforced regret (blue); more often than their Bayesian counterparts, they leave surplus on the table when better options were available. This is by construction; the Bayesian agents are myopically rational and accept all positive-surplus offers. Figure 2 shows that Bayesian agents achieve the highest realized surplus overall, suggesting that this myopic surplus-seeking policy may be well-suited for maximizing value in this environment.
6.3. Qualitative Survey Feedback
Finally, we examine participants’ subjective experiences through post-game survey responses. Despite increasing game complexity, self-reported cognitive exertion, satisfaction, and performance certainty did not differ significantly across the 2-, 3-, and 4-chip conditions. This effort invariance is consistent with the use of fixed-complexity heuristics (such as the balanced ’equal-volume’ trades observed in Figure 3), which may allow the human players to maintain a stable cognitive load regardless of the underlying state-space complexity.
Participants additionally described their trading strategies in words. Thematic analysis (Braun and Clarke, 2006) of these freeform responses revealed three primary themes: profit maximization (cited in 112 responses), collaborative strategy (56), and color optimization (55). [^14] These distributions are visualized in Figure 9. Responses were double-coded if they addressed multiple themes. Profit maximization was the most frequently cited approach (112 responses), with participants describing goals like maximizing their own chip gains. Collaborative and color-focused strategies appeared at similar rates (56 and 55 responses, respectively). Many participants prioritized fairness and mutual benefit, with explanations such as “I aimed to be fair and reasonable with my trades,” or “I kept my offers low to encourage cautious players to engage.” Others took a more aggressive approach, such as “I traded aggressively early on to secure chips before others could deplete the supply.”
These qualitative accounts align with the observed human tendency toward reciprocal exchanges, where participants request a volume of chips similar to the amount offered. These accounts can also be interpreted alongside the self-reported effort invariance. For example, color-based strategies can help players simplify decision-making and manage their cognitive load. Notably, however, none of these self-reported strategies (collaborative, profit-maximizing, or color-optimizing) correlated with actual game performance.
7. Discussion
We briefly discuss behavioral patterns observed across agent types and offer recommendations for deployment and evaluation.
7.1. Empirical Findings
Human behaviors.
Our study sample comprised online strangers engaging in one-shot, mostly anonymous interactions. There was effectively no reputational risk between players—they did not even know whether they were playing against the same humans in the second game. Given this, a traditional game-theoretic model might predict more extractive behaviors prioritizing individual utility maximization (Rubinstein, 1982; Young, 2007). However, the humans offered trades consistent with fairness norms, i.e., the 1:1 ratio of chips offered to chips requested. After the games, themes like coordination and fairness emerged as primary objectives. Such proposals generated less total surplus than Bayesian agents but were more frequently accepted, though humans still exhibited substantial unforced regret, leaving surplus on the table even when better trades were available. In sum, these behaviors could suggest that humans anticipate that their human counterparts will reject unfair offers, and may be more individually rational than overall surplus would suggest.
These findings reveal a significant tension between traditional game-theoretic predictions and observed human behavior. Even in an environment with anonymity and one-shot interactions, specifically designed to minimize social accountability, human behavior remains conspicuously non-extractive. This also suggests that human objectives are not easily reduced to standard reward functions; rather, strategic reasoning appears deeply rooted in social context and background knowledge that non-human agents may fail to capture. In a mixed-agent future, it is unclear whether these human fairness norms will hold against non-human negotiators who may not share such norms.
LLM behaviors.
Despite the novel domain, some GPT-4o-powered agents achieved human-level surplus with almost no additional prompting. They achieved this surplus through offering generally concessionary proposals —giving more chips than they requested—which were readily accepted by other LLM agents. However, this acceptance-maximizing approach led to high rates of forced regret, as agents committed chips early and foreclosed superior opportunities later in the game.
While the underlying causes of this baseline conservative play style remain opaque, it likely arises from three intersecting factors: (i) instruction-tuning on cooperative, information-sharing dialogues where minimizing friction is implicitly rewarded (Wei et al., 2022); (ii) a tendency toward risk-averse, passive responses exacerbated by a lack of outcome-driven feedback during the task (Ouyang et al., 2022); and (iii) a learned aversion to asymmetric outcomes, where generous offers serve as a heuristic to secure rapid coordination (Fisher et al., 2011).
The implications of this default behavior are significant. In non-cooperative settings such as zero-sum negotiations, this overly concessionary tendency could lead LLMs to systematically over-compromise at baseline, in the absence of explicitly prompted behavioral changes. Unless post-training techniques evolve to explicitly account for strategic context, designers may need to intentionally steer these models to prevent such biases from being exploited by more extractive counterparts in competitive environments.
The performance gap between GPT-4o and Gemini 1.5 Pro may also reflect differences in model development cycles. Gemini 1.5 Pro was released approximately three months prior to GPT-4o, and the rapid pace of capability improvements in this period could account for some of the observed differences, though we cannot isolate model architecture from training data or other factors.
Bayesian behaviors.
Despite their trades being frequently rejected, Bayesian agents consistently captured the highest surplus. Their proposals were predominantly no-regret, reflecting strategic precision that held even as game complexity increased. While impressive, this strong performance (in terms of surplus) required us to write a bespoke statistical algorithm specifically aligned to the game’s incentives. We have no reason to believe these behaviors will effectively translate to real-world negotiations, especially in repeated interactions requiring trust and reciprocity. Extending Bayesian agents to more socially compatible trading would require richer models of preferences, fairness, and adaptive behavior, potentially increasing complexity and decreasing robustness.
7.2. Direct Implications
Beyond the empirical results, a central contribution of this work is demonstrating the limits of outcome-based evaluation for agentic negotiation systems. Historically, models are trained and benchmarked using static, aggregate metrics such as total surplus. Had our analysis stopped at this level, one might reasonably conclude that some LLMs are functional substitutes for human negotiators. However, we find substantial heterogeneity in how these LLMs behave relative to their human counterparts, both in their trading behavior and in the opportunities they later forego. While uncovering these procedural discrepancies in this study required novel evaluation methods that go beyond current, static agentic benchmarks, these procedural differences may be critical in a practitioner’s decision to deploy agentic systems in a real-world scenario.
Additionally, if one were considering deployment in this trading game, it is not clear whether LLMs or Bayesian agents are the better choice. For satisficing performance, we find that LLMs can generalize well to novel negotiation domains with minimal context-specific prompt-engineering. Indeed, additional prompting provided no advantage over the baseline. Both LLMs and Bayesian agents resolve negotiations near-instantaneously and are relatively costless at the margin compared to humans. These efficiency advantages make automated agents appealing for high-volume or time-sensitive negotiations, even when beliefs about the other agents are imperfect.
For optimal performance, tradeoffs are more complex. Human agents appear to trade according to “unwritten” fairness rules. Bayesian agents maximize surplus but incur frequent rejections that could prove costly in repeated or reputation-sensitive settings.
Given the current model capabilities and limitations, an optimal implementation may involve a modular hybrid approach combining human values with the speed and reasoning of automated agents. Here, we can imagine LLMs paired with planning or optimization components, such as Bayesian tools for valuation inference, inventory reasoning, or limited lookahead, to improve strategic foresight. Humans as reviewers can then evaluate the final decisions to ensure they accord with the social context.
Concretely, we recommend that both researchers studying human–AI alignment and practitioners building or deploying AI negotiation agents in collective settings—especially those where agents operate alongside humans or act on their behalf—evaluate (i) outcome metrics, such as deal rate and achieved surplus; (ii) procedural metrics, such as strategic behaviors; and (iii) alignment metrics, assessing whether an agent’s strategy reflects the intent, risk tolerance, and normative expectations of the human counterpart.
7.3. Societal Implications
Our findings also have broader societal implications for the deployment of AI agents in human-facing negotiation and coordination settings. In our study, humans consistently proposed fair trades despite no explicit incentives for doing so. One plausible explanation is that participants understood their counterparts to be other humans who value fairness, coordination, and mutual acceptance, and adjusted their behavior accordingly. In this sense, human negotiation behavior appears to be shaped not only by material incentives but also by beliefs about the counterpart’s identity and normative expectations. This is consistent with much of the behavioral game theory literature (Camerer, 2003).
These dynamics raise important questions for increasingly mixed-agent environments, where such beliefs may no longer be well calibrated. If humans come to recognize that they are negotiating with concessionary LLMs, they may rationally adopt more aggressive strategies. Conversely, repeated interaction with highly extractive, optimization-driven agents may undermine human fairness norms altogether, shifting negotiations toward more adversarial and less cooperative equilibria. In both cases, agents that achieve comparable surplus through fundamentally different negotiation styles may produce unexpected and differential outcomes in new settings.
8. Limitations and Future Work
Our study focuses on a single, stylized negotiation game with static valuations and single-shot interactions, conditions that align well with Bayesian agents’ task-specific optimization and may exaggerate their advantage. While humans and LLMs bring general-purpose reasoning to the task, their broader capabilities in dynamic, multi-round, or reputation-driven negotiations remain unexplored. Furthermore, our LLMs used minimal prompting, reflecting realistic deployment but limiting exploration of more strategic behaviors from alternative prompts or fine-tuning.
Beyond performance metrics, our analysis reveals distinct behavioral profiles but does not disentangle whether these arise from reasoning limitations, social norms, or inductive biases inherent to each system. Future work should investigate how humans perceive and trust AI negotiators with different behaviors. Another promising avenue for future work is to allow agents to communicate with natural language. Finally, developing methods to steer LLM agents towards desired procedural norms represents a significant ongoing challenge and opportunity. To support this agenda, we’ve released our game implementation on Deliberate Lab (Qian et al., 2025b), enabling researchers to benchmark new models against the human, LLM, and Bayesian baselines established here as they emerge.
9. Conclusion
As LLMs begin to match human-level outcomes in social decision-making tasks, the central question has shifted from whether they can perform to how they do so—and whether their strategies are robust, predictable, and normatively acceptable.
This study provides a controlled empirical approach for comparing humans, LLMs, and bespoke statistical agents in a dynamic bargaining setting. We organize our findings around two questions: how these agents perform in terms of achieved surplus, and how they behave in the process of getting there. For RQ1 (performance), Bayesian agents capture the most surplus, while GPT-4o-based agents can achieve human-level surplus. For RQ2 (process), we show that similar aggregate outcomes can arise from sharply different negotiation styles: humans tend toward fairness-constrained exchange, while LLMs default to concessionary, acceptance-seeking proposals.
These procedural differences are likely to matter more as LLM agents move beyond stylized games and static interactions into longer-horizon, higher-stakes interactions embedded in ongoing relationships. Agents that concede too readily invite exploitation, while agents that extract too aggressively risk rejection, mistrust, or norm violations. Because identifying these failure modes depends on strategic analysis rather than outcome comparison, static evaluations can be misleading. Responsible deployment, therefore, requires process-sensitive evaluation that makes strategic posture legible, quantifying missed opportunities and checking alignment with human stakeholders’ goals, risk tolerance, and fairness norms.
10. GenAI Usage Disclosure
LLMs were deployed as one of our populations of study. Portions of Figure 1 were generated using GPT-4o. LLM-assisted tools were used for code editing in data analysis notebooks, proofreading, and light copyediting. Generative AI was not used in other parts of the research process, including literature review, substantial data analysis, or writing. The authors retain full responsibility for all content and conclusions presented herein.
Appendix A Algorithms and Proofs
A.1. Empirical Computation of Increasing Game Complexity
Claim.
The decision space increases quadratically as the number of chip colors increases.
Proof.
This is a simple consequence that in games with $k$ different colors of chips, there are $\binom{k}{2}$ possible pairs of chips to transact. Since the distribution of opponents’ valuations is symmetric across chips (except the green numeraire chip), each proposer must consider $O(k^{2})$ different feasible trades. Indeed, if we empirically compute the expected number of myopically rational trades [^15] We define a myopically rational trade to be one where the proposer gains surplus and assigns some probability that at least one opponent would gain surplus by accepting the trade. in the starting configuration of the 2, 3, and 4-chip game, we arrive at 37.1, 120.7, 250.7 respectively. That is, the 3-chip game has approximately $\binom{3}{2}$ as many myopically rational trades in expectation as the 2-chip game, and the 4-chip game has approximately $\binom{4}{2}$ times as many. This approximation becomes more precise as the number of chips increases, since the numeraire’s relative impact is reduced.
A.2. Proof of No Dominant Strategy
Claim.
No agent $i\in\mathbf{I}$ in the game $\mathcal{M}$ has a dominant strategy.
Setup.
Let $\mathbf{I}$ be the set of agents, $\mathbf{G}$ be the set of goods. For each agent $i\in\mathbf{I}$ and good $g\in\mathbf{G}$, there is a valuation $v_{ig}\in\mathbb{R}_{\geq 0}$. The (feasible) allocations are given by $a_{ig}$, the amount of good $g$ allocated to agent $i$.
The welfare of agent $i$ is
$w_{i}\;=\;\sum_{g\in\mathbf{G}}\,v_{ig}\,a_{ig}.$
The total social welfare is
$w\;=\;\sum_{i\in\mathbf{I}}\,w_{i}\;=\;\sum_{i\in\mathbf{I}}\sum_{g\in\mathbf{G}}v_{ig}\,a_{ig}.$
Let each agent $i$ have a strategy space $\mathcal{S}_{i}$. A (pure) strategy $s_{i}\in\mathcal{S}_{i}$ is a complete description of agent $i$’s behavior in the market (e.g., how they propose, reject, accept, etc.). We write $\mathbf{s}_{-i}$ for a strategy profile of all agents other than $i$. Let $u_{i}(s_{i},\mathbf{s}_{-i})$ denote agent $i$’s resulting payoff (or final utility/welfare) under strategy $s_{i}$ against opponents’ strategies $\mathbf{s}_{-i}$. A strategy $s_{i}^{*}$ is called dominant for agent $i$ if, for all $\mathbf{s}_{-i}\in\mathcal{S}_{-i}$,
$u_{i}\bigl(s_{i}^{*},\mathbf{s}_{-i}\bigr)\;\;\geq\;\;u_{i}\bigl(s_{i},\mathbf{s}_{-i}\bigr)\quad\text{for all }s_{i}\in\mathcal{S}_{i}.$
In other words, $s_{i}^{*}$ guarantees the highest possible utility for $i$, irrespective of what other players do.
Proof.
Assume, by way of contradiction, that there exists a dominant strategy $s_{i}^{*}$ for some agent $i$. By the definition of dominance, $s_{i}^{*}$ must satisfy
$u_{i}\bigl(s_{i}^{*},\mathbf{s}_{-i}\bigr)\;\geq\;u_{i}\bigl(s_{i},\mathbf{s}_{-i}\bigr)\quad\text{for all }s_{i}\in\mathcal{S}_{i}\text{ and for all }\mathbf{s}_{-i}\in\mathcal{S}_{-i}.$
Construct two different scenarios of other agents’ strategies and valuations, denoted by $\mathbf{s}_{-i}^{A}$ and $\mathbf{s}_{-i}^{B}$. These scenarios can differ in (i) how other agents value the goods $\{v_{jg}:j\neq i\}$, and (ii) the demands or offers they make (i.e. how $\mathbf{s}_{-i}$ translates into allocations $a_{jg}$).
  • In scenario $A$, suppose the other agents are willing to trade or allocate a certain good $g^{*}$ to agent $i$ only if $i$ offers a high price or concedes in some manner. A specialized alternative strategy $s_{i}^{A}$ (instead of $s_{i}^{*}$) might yield a strictly higher payoff if $i$ trades aggressively for $g^{*}$.
  • In scenario $B$, suppose the others behave differently (e.g., they no longer place much value on $g^{*}$ but highly value some other good). Now, a different specialized strategy $s_{i}^{B}$ might yield a strictly higher payoff (because the trading strategy for $g^{*}$ in scenario $A$ is no longer beneficial here).
Contradiction of dominance. Since $A$ and $B$ are feasible opponent profiles, no single $s_{i}^{*}$ can weakly dominate both tailored responses $s_{i}^{A}$ and $s_{i}^{B}$ across them. In particular,
(1) $\exists\,\mathbf{s}_{-i}^{A}\;:\;$ $u_{i}\!\left(s_{i}^{A},\mathbf{s}_{-i}^{A}\right)\;>\;u_{i}\!\left(s_{i}^{*},\mathbf{s}_{-i}^{A}\right),$
(2) $\exists\,\mathbf{s}_{-i}^{B}\;:\;$ $u_{i}\!\left(s_{i}^{B},\mathbf{s}_{-i}^{B}\right)\;>\;u_{i}\!\left(s_{i}^{*},\mathbf{s}_{-i}^{B}\right).$
This contradicts the definition of $s_{i}^{*}$ as a dominant strategy.
Because the above argument does not rely on any special assumption beyond the existence of at least two different plausible profiles $\mathbf{s}_{-i}^{A}$ and $\mathbf{s}_{-i}^{B}$, we conclude that no strategy for agent $i$ can be dominant. Since $i$ was arbitrary, no agent in the game has a dominant strategy. $\quad\square$
A.3. Multi-Agent Trading Algorithm with Bayesian Learning
Algorithm 1 Multi-agent Trading Algorithm with Bayesian Learning
1:Inputs:
2:   $\mathbf{I}$: set of players
3:   $\mathbf{G}$: set of goods
4:   $\mathbf{A^{0}}$: initial allocation $\{a_{ig}\}$ for all $i\in\mathbf{I},g\in\mathbf{G}$
5:   $\{B_{i}(\mathbf{v}_{-i})\}_{i\in\mathbf{I}}$: each player’s prior belief over other players’ valuations
[…]
7:for all $i\in\mathbf{I}$ do
8:  initialize $B_{i}(\mathbf{v}_{-i})$ $\triangleright$ Each agent’s prior over other agents’ valuations
9:end for
[…]
11:while possible_tradespossible\_trades do
12:  $trades\_found\leftarrow\texttt{false}$
13:  $I_{\text{shuffled}}\leftarrow\text{RandomShuffle}(\mathbf{I})$
14:  for $k=1$ to $|\mathbf{I_{\text{shuffled}}}|$ do
15:    for $l=k+1$ to $|\mathbf{I_{\text{shuffled}}}|$ do
16:     $(i,j)\leftarrow(I_{\text{shuffled}}[k],I_{\text{shuffled}}[l])$
17: $(\mathbf{A^{*}_{ij}},\,\textit{outcome},\,r)\leftarrow\textsc{SolveA}^{*}\Bigl(i,j,\mathbf{G},\{a_{ig},a_{jg}\}_{g\in\mathbf{G}},B_{i},B_{j}\Bigr)$ $\triangleright$ Proposer $i$ finds best trade; $r$ is the responding agent (who accepts/rejects)
18:     if $\textit{outcome}=\textsc{accept}$ then $\triangleright$ Trade executed
19:      $\{a_{ig},\,a_{jg}\}_{g\in\mathbf{G}}\leftarrow\mathbf{A^{*}_{ij}}$
20:      $trades\_found\leftarrow\texttt{true}$
21:     end if
22:     $\textsc{BayesianUpdate}\bigl(i,\,j,\,r,\,\textit{outcome},\,B_{i},\,B_{j},\,\{B_{k}\}_{k\neq i,j}\bigr)$ $\triangleright$ All players update beliefs based on accept/reject
23:    end for
24:  end for
25:  $possible\_trades\leftarrow trades\_found$
26:end while
27:return $\mathbf{A}$
Explanation of key subroutines:
  • $\textsc{SolveA}^{*}(i,j,\dots)$:
    • Proposer’s optimization. Agent $i$ (the proposer) solves
      (3) $\max_{(\mathbf{x},\mathbf{y})}$ $\sum_{\mathbf{v}_{-i}}\mathbf{1}\!\left\{\exists r\neq i:\ \textsc{accept}\bigl(v_{r};\mathbf{x},\mathbf{y}\bigr)\right\}$
      $\qquad\qquad\times\bigl(u_{i}(h_{i}-\mathbf{x}+\mathbf{y})-u_{i}(h_{i})\bigr)\times B_{i}(\mathbf{v}_{-i}).$
      where $\mathbf{x}$ are the chips given up by $i$ and $\mathbf{y}$ are the chips requested, $h_{i}$ denotes $i$’s current holdings, and $B_{i}$ is $i$’s belief over others’ valuations.
    • Receiver’s acceptance criterion. A potential responder $r$ accepts the trade $(\mathbf{x},\mathbf{y})$ myopically if
      $u_{r}\bigl(h_{r}-\mathbf{y}+\mathbf{x};\,v_{r}\bigr)\;>\;u_{r}\bigl(h_{r};\,v_{r}\bigr).$
      The subroutine returns $\mathbf{A^{*}_{ij}}$ (the updated allocation for $i$ and $j$), the outcome (accept or reject), and the identity of the actual responder $r$ (in case multiple receivers are considered, or if $j$ is always the designated responder, $r=j$).
  • $\textsc{BayesianUpdate}\bigl(i,j,r,\textit{outcome},B_{i},B_{j},\{B_{k}\}_{k\neq i,j}\bigr)$:
    • On acceptance: All agents discard any valuation states inconsistent with $r$ accepting $(\mathbf{x},\mathbf{y})$. Specifically,
      $v_{r}\text{ is retained only if }u_{r}(h_{r}-\mathbf{y}+\mathbf{x};v_{r})>u_{r}(h_{r};v_{r}).$
    • On rejection: All agents discard any valuation states for $r$ that would have accepted $(\mathbf{x},\mathbf{y})$, since that contradicts the observed rejection.
    • Then, each agent renormalizes its belief distributions $B_{k}$ to sum to 1 over the remaining (still-plausible) valuations.
This procedure repeats until no further beneficial trades can be found. The Bayesian updates ensure that, over time, agents refine their beliefs about each other’s valuations based on observed accept/reject decisions, thereby learning which trades are more likely to succeed.
Appendix B Human Data Collection
B.1. Implementation on Deliberate Lab
Participant interface.
Upon entering Deliberate Lab through a web link, participants proceed through a multi-stage experiment including informed consent, game instructions, comprehension checks, and payout information. Upon completing the final comprehension check, they wait in a “Lobby” stage for other participants. When three participants are in the lobby, they are sent an invitation to join a live bargaining game (Figure 6). Following the game, there is a post-game survey. For anonymity, we used a Deliberate Lab feature that assigns participants an anonymous animal avatar (e.g., “Bear”) as they join the experiment.
Figure 6. The game interface exactly as it was shown to participants. The top left indicates that they are currently the proposer, offering an interface to propose an offer. A table below displays all players’ chip volumes. To the right, the public ledger shows a history of game rounds, turns, and trade statuses.
The game interface exactly as it was shown to participants. The top left indicates that they are currently the proposer, offering an interface to propose an offer. A table below displays all players’ chip volumes. To the right, the public ledger shows a history of game rounds, turns, and trade statuses.
Figure 7. Experiment dashboard for managing participants. On the left, experimenters can view participants as they enter the experiment, send attention checks, and/or remove participants. On the right, experimenters can monitor the status of a selected participant and assign them to a cohort of three for gameplay.
Experiment dashboard for managing participants. On the left, experimenters can view participants as they enter the experiment, send attention checks, and/or remove participants. On the right, experimenters can monitor the status of a selected participant and assign them to a cohort of three for gameplay.
Bargaining game user interface.
To submit trade proposals, participants used text and selection fields to specify the number and type of chip, respectively. The interface automatically calculated and displayed the projected total chip value if the proposal were to be accepted (Figure 6); it also prevented players from submitting invalid offers by disabling the “submit” button. For accepting or rejecting proposals, the interface calculated and displayed the projected total chip value if the player were to accept the proposal. It also disabled acceptance for offers the player could not fulfill (Figure 8, bottom). All participants were shown a table of all players’ chip quantities and a log of actions (i.e., when trades were proposed, accepted, and rejected, and when each round and turn of the game changed). Turn order was determined randomly.
Experimenter interface.
To conduct experiments, the experimenter used Deliberate Lab’s dashboard (Figure 7) to monitor participants as they progressed through the experiment’s information and comprehension check stages, then manually transferred participants into groups of three to play the bargaining game once the player was ready to be transferred. The experimenter was also able to send attention checks to participants (e.g., if several minutes elapsed without any visible progress) and remove them from the experiment (e.g., if attention checks went unanswered).
Figure 8. Accepting or rejecting a trade. The bottom image shows a disabled “accept” option.
Accepting or rejecting a trade. The bottom image shows a disabled “accept” option.
Appendix C Post-Game Participant Survey
Game variation Competitive Certainty Satisfaction Exertion
2-chip 5.57 (0.34) 5.57 (0.33) 6.68 (0.31) 7.11 (0.24)
3-chip 5.32 (0.34) 5.92 (0.33) 7.07 (0.30) 7.10 (0.25)
4-chip 5.75 (0.34) 6.04 (0.33) 7.46 (0.26) 7.46 (0.19)
Table 1. Summary statistics of self-assessment Likert variables (mean and standard error); full survey questions in Table 2. There was no significant change in certainty or exertion as game complexity increased.
The survey questions are provided in Table 2. Responses to Likert-scale questions are aggregated in Table 1. Qualitative thematic analysis is provided in Figure 9.
Table 2. Post-game survey questions.
Survey question Response format
1 Please describe your strategy in the game in a few sentences. Freeform text
2 Please describe any experiences or background that may have influenced your performance in the game. Freeform text
3 On a scale from 1 to 10, how would you rate your trading strategy, where 1 is highly competitive (focused mainly on your own gains) and 10 is highly collaborative (focused on other players’ potential gains and mutual benefits) 1 through 10 scale; 1: “highly competitive”, 10: “highly collaborative”
4 On a scale from 1 to 10, how certain are you that you made the best possible trades during the experiment? 1 through 10 scale; 1: “Not at all confident”, 10: “Very confident”
5 On a scale from 1 to 10, how satisfied are you with your final trading outcomes? 1 through 10 scale; 1: “Not at all satisfied”, 10: “Very satisfied”
6 On a scale of 1 to 10, how would you rate the mental effort you put into today’s trading games, where 1 means ’I barely engaged in any thinking’ and 10 means ’I put in significant mental effort’? 1 through 10 scale; 1: “No mental exertion at all”, 10: “Very high mental exertion”
7 Please provide any additional context on your answers above. Freeform text
8 Please help us to improve this experiment. How was your experience today? Were there any elements of the instructions or gameplay that you found confusing? Freeform text
Figure 9. Distribution of self-reported strategies across game complexity (thematic analysis). Participants described their trading approach as focused on profit maximization, collaborative play, or color optimization. Profit-focused strategies were most common (112 responses), while collaborative (56) and color-focused (55) strategies appeared at similar rates. The distribution across 2-, 3-, and 4-chip games was relatively consistent within each strategy type.
Distribution of self-reported strategies across game complexity (thematic analysis). Participants described their trading approach as focused on profit maximization, collaborative play, or color optimization. Profit-focused strategies were most common (112 responses), while collaborative (56) and color-focused (55) strategies appeared at similar rates. The distribution across 2-, 3-, and 4-chip games was relatively consistent within each strategy type.
Appendix D Additional Clarification on Strategic Regret Classifications
Figure 10. Regret behavior visualization: The top-left panel shows an overview trajectory from Figure 4. The top-right panel illustrates a “No Regret” scenario for P2 as a proposer. The bottom-left panel depicts “Forced Regret” for P3 as an acceptor. The bottom-right panel shows “Unforced Regret” for P2 as a decliner at Turn 8. Solid lines represent actual player paths, while dashed lines indicate counterfactual paths.
[…]
Here, we utilized counterfactual analysis to visually dissect the three kinds of regret behavior previously discussed in Section 6.2.2, as illustrated in Figure 10. In these visualizations, the player’s actual path of total chip value is depicted by a solid colored line, while the counterfactual path—representing the outcome had a different decision been made—is indicated by a gray dashed line.
Counterfactual definition: The counterfactual path is constructed by simulating an alternative decision at a critical juncture (e.g., accepting a declined offer, declining an accepted offer, or making a different proposal) and projecting the subsequent value trajectory. This allows for a direct comparison between the actual outcome and what could have been.
No regret scenarios. This scenario is demonstrated by Player 2 (P2) in the role of a proposer, as shown in the upper-right panel of Figure 10, focusing on Turn 9. Player 2 makes a proposal that is accepted, leading to an actual chip value of $15.3$. The counterfactual path, representing a scenario where P2 might not have made this proposal or made a less optimal one, results in a lower value of 14.9. In this “No Regret” instance, the player’s actual total value path is superior to or equal to the evaluated counterfactual path, indicating that the decision made was optimal given the available information and subsequent events.
Forced regret scenarios. The lower-left panel of Figure 10 illustrates “Forced Regret” experienced by Player 3 (P3) as an acceptor from Turn 1 to Turn 2. P3 initially accepts an offer from P1 at Turn 1, resulting in an actual value of 10.8. However, a more lucrative offer from P2 becomes available in Turn 2. Due to the commitment made in Turn 1 (e.g., changed chip inventory), P3 is unable to capitalize on this subsequent, better opportunity. The counterfactual path shows that had P3 declined P1’s initial offer, they could have potentially accepted P2’s offer, leading to a higher value of 11.0.
Unforced regret scenarios. The lower-right panel of Figure 10 demonstrates Player 2 (P2) in the role of a decliner from Turn 8 to 9. P2 is presented with a profitable trade offer but chooses to decline it. P2’s actual chip value remains at 14.9 after this decision. The counterfactual path, however, demonstrates that had P2 accepted this profitable offer, their chip value would have increased to 15.0. This highlights an avoidable error: a superior option was available and actionable, but the agent failed to select it, leading to a sub-optimal outcome. Unlike forced regret, no prior decision prevented P2 from taking the better option; the regret stems directly from the choice made at that specific juncture.
Appendix E Surplus Values and Significance
Surplus values are provided in Table 3 and Table 4.
Population 2-chip 3-chip 4-chip
Human 0.60 (0.06) 0.59 (0.04) 0.54 (0.03)
GPT-4o 0.69 (0.03) 0.62 (0.03) 0.54 (0.02)
GPT-4o refined 0.68 (0.04) 0.64 (0.03) 0.58 (0.02)
Gemini 0.42 (0.04) 0.42 (0.03) 0.37 (0.02)
Gemini refined 0.45 (0.05) 0.43 (0.03) 0.33 (0.02)
Bayesian agent 0.74 (0.04) 0.80 (0.03) 0.73 (0.02)
Table 3. This table provides the means and standard errors for the average ratio of observed total surplus to optimal surplus achieved across agent-game combinations. It corresponds to the numbers in Figure 2.
Human GPT-4o GPT-4o refined Gemini Gemini refined Bayesian agent
2-chip
Human 0.491 ↑ 0.597 ↑ 0.263 ↓ 0.066 ↓ 0.654 ↑
GPT-4o 0.302 ↓ $p<0.001$ $p<0.001$ 0.141 ↑
GPT-4o refined $p<0.001$ $p<0.001$ 0.414 ↑
Gemini 0.359 ↑ $p<0.001$
Gemini refined $p<0.001$
Bayesian agent
3-chip
Human 0.525 ↑ 0.021 ↑ 0.001 ↓ $p<0.001$ $p<0.001$
GPT-4o 0.164 ↑ $p<0.001$ 0.004 ↓ $p<0.001$
GPT-4o refined $p<0.001$ $p<0.001$ 0.001 ↑
Gemini 0.799 ↑ $p<0.001$
Gemini refined $p<0.001$
Bayesian agent
4-chip
Human 0.601 0.258 ↑ $p<0.001$ $p<0.001$ $p<0.001$
GPT-4o 0.270 ↑ $p<0.001$ $p<0.001$ $p<0.001$
GPT-4o refined $p<0.001$ $p<0.001$ $p<0.001$
Gemini 0.003 ↓ $p<0.001$
Gemini refined $p<0.001$
Bayesian agent
Table 4. Two-sided t-test p-values comparing surplus values across populations (144 games). Arrows indicate whether the column’s mean surplus was higher (↑) or lower (↓) than the row’s.
Appendix F Additional Data Analysis
F.1. Learning Effects
In Figure 11, we investigate learning effects for humans in the 2, 3, and 4-chip variants of the negotiation game by comparing the final surplus gain across the two stages of play. We see that while there is a small increase in performance in each game, it is not statistically significant. We also cannot determine how the complexity of the game relates to learning effects, as the 3-chip game seems to have the smallest effect.
Figure 11. Between stage (game) comparisons of human group surplus by game variant.
Between stage (game) comparisons of human group surplus by game variant.
F.2. Supplementary Trading Pattern Visualizations
Summary statistics of trading patterns are provided in Table 5, and average final surplus information in Table 3. The 2-chip visualization is shown in Figure 12 and 4-chip visualization in Figure 13. They exhibit similar behaviors as discussed in Section 6.2.1.
Figure 12. This figure is identical to Figure 3 in the main text, but provides the trading patterns for the 2-chip game.
[…]
Figure 13. This figure is identical to Figure 3 in the main text, but provides the trading patterns for the 4-chip game.
[…]
Game Player Acceptance Surplus Mean (SD) Ratio Mean (SD) Median [Surplus, Ratio]
2-chip Human Accepted 0.611 (1.031) 1.147 (0.652) [0.500, 1.000]
Rejected 1.088 (1.313) 0.735 (0.368) [0.800, 0.667]
GPT-4o Accepted 0.148 (0.617) 1.711 (0.976) [0.100, 1.500]
Rejected 0.414 (1.031) 0.715 (0.393) [0.100, 0.667]
Gemini-1.5 Pro Accepted -0.305 (0.913) 1.508 (0.721) [-0.100, 1.500]
Rejected 0.230 (1.317) 1.048 (0.475) [0.100, 1.000]
Bayesian Accepted 2.245 (1.252) 0.867 (0.315) [2.300, 0.900]
Rejected 2.058 (1.904) 0.636 (0.372) [1.500, 0.526]
3-chip Human Accepted 1.211 (1.317) 1.039 (0.387) [1.100, 1.000]
Rejected 1.830 (1.803) 0.840 (0.325) [1.300, 1.000]
GPT-4o Accepted 0.437 (0.838) 1.765 (0.945) [0.200, 1.600]
Rejected 0.345 (0.963) 0.918 (0.501) [0.200, 0.800]
Gemini-1.5 Pro Accepted 0.329 (0.899) 1.394 (0.610) [0.400, 1.500]
Rejected 0.528 (1.819) 1.075 (0.497) [0.600, 1.000]
Bayesian Accepted 3.173 (1.853) 0.944 (0.274) [3.200, 0.900]
Rejected 3.385 (2.725) 0.783 (0.355) [2.800, 0.700]
4-chip Human Accepted 1.331 (1.214) 1.173 (0.562) [1.200, 1.000]
Rejected 1.757 (1.438) 0.863 (0.288) [1.500, 1.000]
GPT-4o Accepted 0.369 (0.929) 1.745 (0.831) [0.200, 1.600]
Rejected 0.534 (0.981) 1.262 (0.518) [0.300, 1.250]
Gemini-1.5 Pro Accepted 0.403 (0.858) 1.438 (0.540) [0.400, 1.500]
Rejected 0.650 (1.187) 1.022 (0.377) [0.600, 1.000]
Bayesian Accepted 3.536 (1.793) 0.965 (0.251) [3.400, 0.900]
Rejected 4.251 (2.824) 0.738 (0.303) [3.500, 0.700]
Table 5. Summary statistics of the trading patterns in Figures 3, 12, and 13.
Appendix G Aside on Smaller Model Performance
We tested smaller models from both the Gemini and GPT families—Gemini-2.5-Flash and GPT-o4-mini—and found that their overall performance falls short of both human participants and larger models such as GPT-4o, particularly in terms of surplus generation. Even with prompt refinement, the average surplus remains substantially lower.
Procedurally, these small models also behave quite differently from their larger counterparts. As shown in Figure 14, proposals by Gemini-2.5-Flash and GPT-o4-mini are heavily concentrated around a narrow region: nearly zero net surplus gain and a 1:1 trade ratio. Roughly 90% of offers consist of exchanging one chip for exactly one of another type (e.g., 1 red for 1 green). Rarely do small LLMs propose trades that would result in a loss. However, they also rarely propose trades that would benefit them substantially (e.g., a net surplus of >2>2). Prompt refinement slightly broadens the distribution but fails to shift the models away from this dominant 1:1 pattern.
Figure 14. This figure is identical to Figure 3 in the main text, but provides the trading patterns for the smaller models
[…]
This low-variance, risk-averse behavior is characteristic of the “lazy generation” effect reported in small language models (Lambert and Calandra, 2023), and reflects the narrow, template-driven action spaces observed in recent bargaining benchmarks for sub-10B parameter agents (Xia et al., 2024).
Appendix H LLM Simulation Prompts
H.1. Rule Explanation
How the game works:
The game consists of 3 rounds of trading. During each round, each player will have a turn to propose 1 trade. These turns are pre-determined in a random order, and the order stays the same in each round.
Trade proposals:
To propose a trade, a player must:
1. Request a certain quantity of chips of a single color from any other player to get.
2. Specify a certain quantity of chips of a different color to give in return.
Trade rules:
Players cannot offer more chips than they currently hold. For example, if you only have 5 red chips, you cannot offer 6 red chips.
Players cannot trade chips of the same color. For example, you cannot trade red chips for red chips.
Trade completion:
When an offer is presented, all other active participants get a chance to accept or decline. Note: Active participants are those not currently making the offer.
Participants make their decisions simultaneously and privately. The participant who receives the offer is not dependent on who accepts the trade first. Some possible outcomes:
If no one accepts, the trade does not happen, and the turn ends.
If multiple participants accept, one accepting participant is chosen at random to complete the trade with the offering participant. This means that participants cannot choose who they trade with.
If only one participant accepts, the trade will happen.
Key points to remember:
In each round, each player gets to propose one trade and respond to other players trades
You can only propose trades between different colored chips, and cannot offer to give a chip amount that you do not have
When multiple players accept a trade, the trading partner is randomly selected
H.2. Proposer’s Prompt
You are {{name}}.
Your valuations of the different types of chips are: {{preference_description}}.
You now have the following amounts of each chip: {{item}}.
The conversation history so far is {{history}}.
REMEMBER, to propose a trade, you must:
Request a certain quantity of chips of a single color from any other player to get
Specify a certain quantity of chips of a different color to give in return
REMEMBER you have the following amounts of each chip: {{item}}.
Your goal is to make as much money as possible. The trades, you choose to make to accomplish this, are up to you.
As a part of making money you must be rational - do not propose a trade in which you lose money. The value of a trade to you is the difference between the total value of chips you receive (quantity x your valuation) minus the total value of chips you give up (quantity x your valuation). Only propose trades that give you positive value.
In short, your trades should be both incentive compatible and individually rational.
Your response must use these EXACT tags below. The response should include nothing else besides the tags, your trade offer, and your reasoning. The text between tags should be concise.
<REASONING>
[Provide your concise reasoning in a few sentences, e.g. To gain more surplus, I want more xxx chips]
</REASONING>
<CHECK>
[check if you have sufficient chips to trade. If you have n green chips, you can at most give n green chips. If you dont want to trade, you can ask for a large amount of chips that no one can afford]
</CHECK>
<GET_COLOR> Color, e.g. red</GET_COLOR>
<GET_QUANTITY> quantity, e.g. n </GET_QUANTITY>
<GIVE_COLOR> Color, e.g. red</GIVE_COLOR>
<GIVE_QUANTITY> quantity, e.g. n </GIVE_QUANTITY>
H.3. Receiver’s Prompt
You are {{name}}.
Your valuations of the different types of chips are: {{preference_description}}.
You now have the following amounts of each chip: {{item}}.
The conversation history so far is {{history}}.
You have an offer. {{proposer}} is offering to give {{give}} and get {{get}} in return.
If you make this trade, your total wealth will change by: {{delta_surplus}}
Now, you need to decide whether to accept or decline.
Your response must use these EXACT tags below. The response should include nothing else besides the tags, your choice to accept or decline, and your reasoning. The text between tags should be concise.
<REASONING>
[Provide your concise reasoning in a few sentences.]
</REASONING>
<CHOICE>Yes or No </CHOICE>
H.4. Refined Proposing Prompt
H.4.1. Multiple proposal generation prompt
You are {{name}}.
Your valuations of the different types of chips are: {{preference_description}}.
You now have the following amounts of each chip: {{item}}.
The conversation history so far is {{history}}.
REMEMBER, to propose a trade, you must:
Request a certain quantity of chips of a single color from any other player to get
Specify a certain quantity of chips of a different color to give in return
REMEMBER you have the following amounts of each chip: {{item}}.
Your goal is to make as much money as possible. The trades, you choose to make to accomplish this, are up to you.
As a part of making money you must be rational - do not propose a trade in which you lose money. The value of a trade to you is the difference between the total value of chips you receive (quantity $\times$$ your valuation) minus the total value of chips you give up (quantity $\times$$ your valuation). Only propose trades that give you positive value.
In short, your trades should be both incentive compatible and individually rational.
You can trade as many chips as you want in a single turn, assuming you have that many. Do not feel constrained to only trade a single chip at a time.
Propose 3 different good trade ideas, so a next step can decide on the best trade of the ones you propose here.
Your response must use these EXACT tags below. The response should include nothing else besides the tags, your trade offer, and your reasoning. The text between tags should be concise.
Repeat the below tags once for each of the trade ideas you propose.
<REASONING>
[Provide your concise reasoning in a few sentences, e.g. To gain more surplus, I want more xxx chips]
</REASONING>
<CHECK>
[check if you have sufficient chips to trade. If you have n green chips, you can at most give n green chips. If you dont want to trade, you can ask for a large amount of chips that no one can afford]
</CHECK>
<GET_COLOR> Color, e.g. red</GET_COLOR>
<GET_QUANTITY> quantity, e.g. n </GET_QUANTITY>
<GIVE_COLOR> Color, e.g. red</GIVE_COLOR>
<GIVE_QUANTITY> quantity, e.g. n </GIVE_QUANTITY>
H.5. Choosing from Multiple Possible Proposals Prompt
You are {{name}}.
Your valuations of the different types of chips are: {{preference_description}}.
You now have the following amounts of each chip: {{item}}.
The conversation history so far is {{history}}.
REMEMBER, to propose a trade, you must:
Request a certain quantity of chips of a single color from any other player to get
Specify a certain quantity of chips of a different color to give in return
REMEMBER you have the following amounts of each chip: {{item}}.
Your goal is to make as much money as possible. The trades, you choose to make to accomplish this, are up to you.
As a part of making money you must be rational - do not propose a trade in which you lose money. The value of a trade to you is the difference between the total value of chips you receive (quantity $\times$$ your valuation) minus the total value of chips you give up (quantity $\times$$ your valuation). Only propose trades that give you positive value.
In short, your trades should be both incentive compatible and individually rational.
You can trade as many chips as you want in a single turn, assuming you have that many. Do not feel constrained to only trade a single chip at a time.
Below are three trade ideas you have proposed. Please pick the best trade proposal from the ones below that you will propose to the group.
Do not change anything about the trade you have selected from the ideas below.
Your response must use these EXACT tags below. The response should include nothing else besides the tags and content of your selected trade offer including its reasoning.
<REASONING>
[Provide your concise reasoning in a few sentences, e.g. To gain more surplus, I want more xxx chips]
</REASONING>
<CHECK>
[check if you have sufficient chips to trade. If you have n green chips, you can at most give n green chips. If you dont want to trade, you can ask for a large amount of chips that no one can afford]
</CHECK>
<GET_COLOR> Color, e.g. red</GET_COLOR>
<GET_QUANTITY> quantity, e.g. n </GET_QUANTITY>
<GIVE_COLOR> Color, e.g. red</GIVE_COLOR>
<GIVE_QUANTITY> quantity, e.g. n </GIVE_QUANTITY>
Proposed trade ideas to choose from:
{{proposed}}